### v1

# Balanced Mixed-Type Tabular Data Synthesis with Diffusion Models

Abstract
Diffusion models have emerged as a robust framework for various generative tasks, such as image and audio synthesis, and have also demonstrated a remarkable ability to generate mixed-type tabular data comprising both continuous and discrete variables. However, current approaches to training diffusion models on mixed-type tabular data tend to inherit the imbalanced distributions of features present in the training dataset, which can result in biased sampling. In this research, we introduce a fair diffusion model designed to generate balanced data on sensitive attributes. We present empirical evidence demonstrating that our method effectively mitigates the class imbalance in training data while maintaining the quality of the generated samples. Furthermore, we provide evidence that our approach outperforms existing methods for synthesizing tabular data in terms of performance and fairness.
Machine Learning, Diffusion Models
1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) are a class of energy-based generative models utilizing a stochastic process to transform desired data samples from noise iteratively and have demonstrated remarkable performance in generative modeling for various tasks, especially text-to-image synthesis. Researchers use diffusion frameworks to model images paired with text descriptions and show promising results (Nichol et al., 2021; Saharia et al., 2022). Advancing further from diffusion models in pixel space, latent diffusion frameworks optimize image representation in a low-dimensional space, thereby enhancing synthetic image quality and efficiency (Rombach et al., 2022; Ramesh et al., 2022). The impressive results of diffusion models in image synthesis extended to the generation of sequential data such as audio and time series (Kong et al., 2020; Rasul et al., 2021; Li et al., 2022). This advancement has gradually overshadowed generative adversarial networks (previous research spot in generative modeling), as diffusion models consistently demonstrate superior performance across diverse tasks (Dhariwal & Nichol, 2021; Stypułkowski et al., 2023; Mazé & Ahmed, 2023).
Researchers have recently been motivated by the versatility of diffusion models to explore the potential in synthesizing tabular data (Kotelnikov et al., 2023; He et al., 2023). Tabular data often involves sensitive user information, which raises significant privacy concerns when used for model training. Moreover, such datasets suffer from limitations in size and imbalanced class distribution (Jesus et al., 2022), which leads to challenges in training robust and fair machine learning models. To address these issues, there is a growing demand for effective tabular data synthesis. However, generative modeling of tabular data is challenging due to the combination of discrete and continuous features and their varying value distributions. Interpolation-based techniques, such as SMOTE (Chawla et al., 2002), are simple but effective and challenging to beat. Researchers have used adversarial generative networks and variational autoencoders to generate tabular data with promising results (Ma et al., 2020; Xu et al., 2019). Research on diffusion models for tabular data is ongoing, and the primary challenge is to handle continuous and discrete features simultaneously. In (Austin et al., 2021), the authors suggested using diffusion frameworks for discrete features with Markov transition matrices. Recent developments such as TabDDPM (Kotelnikov et al., 2023), which combines Markov transition functions (called diffusion kernels) for continuous and discrete features, mark significant progress in this field and underscore the potential of diffusion models as a comprehensive solution for tabular data synthesis.
Despite the success of diffusion models in various generative tasks, ethical concerns regarding potential bias and unsafe content in the synthetic data persist. Generative models are inherently data-driven, thereby propagating these biases into the synthetic data. Studies such as (Friedrich et al., 2023; Schramowski et al., 2023) have shown that latent diffusion models, particularly in text-to-image synthesis, can create images biased toward certain privileged groups and may contain inappropriate elements, such as nudity and violence. In response, they propose methods to apply hazardous or sensitive guidance in latent space and then eliminate bias and offensive elements in the synthetic images. In addition to text and images, tabular datasets frequently exhibit notable bias, often stemming from issues related to the disproportionate representation of sensitive groups. For instance, some datasets exhibit significant imbalances in demographic groups regarding sensitive attributes such as sex and race (Le Quy et al., 2022). However, in our review of the literature, research on the safety and fairness of diffusion models in areas beyond text-to-image synthesis is limited. Although diffusion models have shown remarkable performance in tabular data synthesis (Sattarov et al., 2023; He et al., 2023; Kotelnikov et al., 2023), they are prone to learning and replicating these biases in training data. To address this issue, we propose a novel diffusion framework to accurately model distributions of mixed-type tabular data, with a specific focus on outcomes and sensitive attributes. We employ balanced sampling techniques to reduce disparities in group representation within the synthetic data. In summary, our main contributions are as follows:
  1. 1.
    We introduce a diffusion-based framework tailored to learn distributions of mixed-type tabular data given multiple attributes.
  2. 2.
    We generate tabular data that is balanced for a predefined set of sensitive attributes, addressing inherent biases in the data.
  3. 3.
    Our model surpasses existing models in specific areas of performance and fairness, as evidenced through comprehensive evaluations.
The remainder of this paper is organized as follows: Section 2 provides an overview of fairness-aware machine learning and diffusion models. Section 3 provides an overview of mixed-type modeling with diffusion models. In Section 4, we detail our approach to multivariate conditioning and balanced sampling. In Section 5, we illustrate the experiments and present the results. Finally, in Section 6, we conclude our findings and discuss their implications.
2 Related Work
The concept of bias in machine learning that involves assigning significance to specific features to enhance overall generalization is essential for the effective performance of models (Mehrabi et al., 2021). However, it is crucial to acknowledge that bias in machine learning can also have negative implications. Negative bias is an inaccurate assumption made by machine learning algorithms that reflects systematic or historical prejudices against certain groups of people (Zanna et al., 2022). Decisions derived from such biased algorithms can lead to adverse effects, particularly impacting specific social groups defined by factors such as sex, age, disability, race, and more.
This concept of bias and fairness in machine learning has been widely studied, the aim being to mitigate algorithmic bias of sensitive attributes from machine learning models (Mehrabi et al., 2021; Pessach & Shmueli, 2022). Many methods for bias mitigation in classification tasks have been proposed, and they can be classified into three categories: pre-processing, in-processing, and post-processing methods. Pre-processing methods (Kamiran & Calders, 2012; Calmon et al., 2017) transform the collected data to remove discrimination and train machine learning models on discrimination-free datasets. In-processing methods (Zhang et al., 2018; Agarwal et al., 2018; Kamishima et al., 2011; Zafar et al., 2017; Kamishima et al., 2012) regulate machine learning algorithms by incorporating fairness constraints or regularization terms into the objective functions. Lastly, post-processing methods (Hardt et al., 2016; Pleiss et al., 2017), implemented after training machine learning models on collected datasets, directly override the predicted labels to improve fairness. Fairness-aware generative models can be classified as pre-processing methods. In one of such methods proposed to generate fair data, (Xu et al., 2018) introduced fairness-aware adversarial generative networks that employ a fair discriminator to maintain equality in the joint probability of features and labels conditioned on subgroups within sensitive attributes. However, this approach does not effectively tackle the imbalance in sensitive classes within synthetic data, and diffusion-based tabular data synthesis methods such as TabDDPM (Kotelnikov et al., 2023) excel in producing synthetic data of superior quality.
With impressive capabilities in generative modeling, diffusion models are a class of deep generative models that utilize a Markov process that starts from noise and ends with the target data distribution. Although diffusion models are widely studied, the research on fairness and safety of diffusion models is limited, and existing works (Schramowski et al., 2023; Friedrich et al., 2023) mainly focus on text-to-image synthesis tasks. Researchers in (Friedrich et al., 2023) utilize sensitive content, such as sex and race, within text prompts to guide training diffusion models. Subsequently, during the sampling phase, a uniform distribution of sensitive attributes is applied to generate images that are fair and do not exhibit preference towards privileged groups. Similarly, (Schramowski et al., 2023) uses unsafe text, such as nudity and violence, to guide diffusion models and remove unsafe content in the sampling phase. In addition, this method applies some tricks to remove unsafe content from synthetic images while keeping changes minimal. However, these fairness-aware diffusion models have not been adapted to tabular data synthesis, even though bias commonly exists in tabular datasets (Le Quy et al., 2022). This paper studies how to model mixed-type tabular data conditioning on labels and multiple sensitive attributes in latent space. Our method can generate balanced tabular data considering multiple sensitive attributes and subsequently achieve improved fairness scores without compromising the quality of the synthetic data.
3 Background
Diffusion models, taking inspiration from thermodynamics, are probabilistic generative models that operate under the Markov assumption. Diffusion models involve two major components: a forward process and a reverse process. The forward process gradually transforms the given data $\mathbf{x}_{0}$ into noise $\mathbf{x}_{T}$ passing through $T$ steps of the diffusion kernel $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$. In the reverse process, $\mathbf{x}_{T}$ is restored to the original data $\mathbf{x}_{0}$ by $T$ steps of posterior estimator $p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ with trainable parameters $\mathbf{\theta}$. In this section, we explain the Gaussian diffusion kernel for continuous features and the multinomial diffusion kernel for discrete features. Furthermore, we show how to optimize the parameters $\mathbf{\theta}$ of the posterior estimator considering mixed-type data with both continuous and discrete features.
3.1 Gaussian Diffusion Kernel
The Gaussian diffusion kernel is used in the forward process for continuous features. After a series of Gaussian diffusion kernels with $T$ timesteps, $\mathbf{x}_{T}$ follows the standard Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$. Specifically, with the values of $\beta_{t}$ predefined according to a schedule such as linear, cosine, etc. (Chen, 2023), the Gaussian diffusion kernel $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ can be written as:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t}|\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I})$
With the starting point $\mathbf{x}_{0}$ known, the posterior distribution in the forward process can be calculated analytically as follows,
$q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t-1}|\tilde{\mathbf{\mu}}_{t}(\mathbf{x}_{t},\mathbf{x}_{0}),\tilde{\beta}_{t}\mathbf{I})$
$\tilde{\mathbf{\mu}}_{t}(\mathbf{x}_{t},\mathbf{x}_{0})=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\mathbf{\epsilon})\$ (1)
$\tilde{\beta}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$. In the reverse process, a posterior estimator with parameters $\mathbf{\theta}$ is used to approximate $p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ at each timestep $t$ given $\mathbf{x}_{t}$.
$p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1}|\hat{\mathbf{\mu}}(\mathbf{x}_{t}),\hat{\mathbf{\Sigma}}(\mathbf{x}_{t}))$
Researchers in (Ho et al., 2020) set the estimated covariance matrix $\hat{\mathbf{\Sigma}}(\mathbf{x}_{t})$ as diagonal matrix with constant values $\beta_{t}$ or $\tilde{\beta}_{t}$, and calculate the estimated mean $\hat{\mathbf{\mu}}(\mathbf{x}_{t})$ as follows,
$\hat{\mathbf{\mu}}(\mathbf{x}_{t})=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\mathbf{\epsilon}}(\mathbf{x}_{t}))$ (2)
where $\hat{\mathbf{\epsilon}}(\mathbf{x}_{t})$ is the estimated noise at time $t$.
3.2 Multinomial Diffusion Kernel
The multinomial diffusion kernel is applied in the forward process for discrete features with the same noise schedule $\beta_{t}$ used by the Gaussian diffusion kernel. For discrete features, in the final timestep $T$, $\mathbf{x}_{T}$ follows the discrete uniform distribution $\operatorname{Uniform}(K)$, where $K$ is the number of categories. The multinomial diffusion kernel can be written as:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\operatorname{Cat}(\mathbf{x}_{t}|(1-\beta_{t})\mathbf{x}_{t-1}+\beta_{t}/K)$
Similarly to the continuous case, the posterior distribution of the multinomial diffusion kernel has an analytical solution.
$q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})=\mathrm{Cat}(\mathbf{x}_{t-1}|\tilde{\mathbf{\theta}}/\sum_{k=1}^{K}\tilde{\theta}_{k})$
$\tilde{\mathbf{\theta}}=[\alpha_{t}\mathbf{x}_{t}+(1-\alpha_{t})/K]\odot[\bar{\alpha}_{t-1}\mathbf{x}_{0}+(1-\bar{\alpha}_{t-1})/K]$
The posterior estimator for discrete features is preferably parameterized by the estimated starting point $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$ as suggested by (Hoogeboom et al., 2021).
$p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))$ (3)
3.3 Model Fitting
The parameters ${\mathbf{\theta}}$ of the posterior estimator are learned by minimizing the variational lower bound,
$\mathcal{L}(\mathbf{x}_{0})=\mathbb{E}_{q(\mathbf{x}_{0})}[D(q(\mathbf{x}_{T}|\mathbf{x}_{0})\|p(\mathbf{x}_{T}))-\log p_{\mathbf{\theta}}(\mathbf{x}_{0}|\mathbf{x}_{1})$ (4)
$+\sum_{t=2}^{T}D(q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})\|p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t}))]$
where Kullback–Leibler divergence is denoted by $D$. The variational lower bound can be simplified for continuous variables as the mean squared error between true noise $\mathbf{\epsilon}$ in Equation 1 and estimated noise $\hat{\mathbf{\epsilon}}$ in Equation 2, denoted by $\mathcal{L}_{\text{G}}$.
$\mathcal{L}_{\text{G}}=\mathbb{E}_{q(\mathbf{x}_{0})}[\|\mathbf{\epsilon}-\hat{\mathbf{\epsilon}}({\mathbf{x}_{t}})\|^{2}]$
For mixed-type data containing continuous features and $C$ categorical features, the total loss $\mathcal{L}_{\text{T}}$ can be expressed as the summation of the Gaussian diffusion loss term $\mathcal{L}_{\text{G}}$ and the average of the multinomial diffusion loss terms in Equation 4 for all categorical features:
$\mathcal{L}_{\text{T}}=\mathcal{L}_{\text{G}}+\frac{\sum_{i\leq C}\mathcal{L}^{(i)}}{C}$
To train latent diffusion models, let $\mathbf{z}_{t}$ be the latent representation of $\mathbf{x}_{t}$, the estimated noise $\hat{\mathbf{\epsilon}}(\mathbf{x}_{t})$ for continuous features in Equation 2 can be reformulated as $\hat{\mathbf{\epsilon}}(\mathbf{z}_{t})$. The estimated original input $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$ in Equation 3 can be reformulated as $\hat{\mathbf{x}}_{0}(\mathbf{z}_{t})$.
4 Methods
As explained in Section 1, generative diffusion models for tabular data are prone to learning the sensitive group size disparity in training data. In this section, we describe the process of synthesizing balanced mixed-type tabular data considering sensitive attributes. We first introduce conditional generation given multivariate guidance while preserving the synthesis quality. Then, we explain the architecture of the backbone deep neural network used as the posterior estimator in our method. Lastly, we explain the procedure for sampling balanced data with consideration for sensitive attributes. The high-level structure of our method is shown in Figure 1.
Figure 1: The diagram of our model architecture. In the forward process, the input data point $\mathbf{x}$ is pre-processed into numerical part $\mathbf{x}_{\text{num}}$ and categorical part $\mathbf{x}_{\text{cat}}$, and then passing through $T$ steps of the diffusion kernel to get $\mathbf{x}_{1},\cdots,\mathbf{x}_{T}$. $\mathbf{z}_{1},\cdots,\mathbf{z}_{T}$ is the latent representation of $\mathbf{x}_{1},\cdots,\mathbf{x}_{T}$. In the reverse process, the posterior estimator iteratively denoises noisy input $\mathbf{z}_{T}$ conditioning on an outcome $\mathbf{c}$ and $N$ sensitive attributes $\mathbf{s}^{(1)},\cdots,\mathbf{s}^{(N)}$. The estimated data point is $\hat{\mathbf{x}}$.
4.1 Multivariate Latent Guidance
Conditional data generation is important in many tasks like text-to-image synthesis. To achieve conditional data generation with diffusion models, classifier-free guidance (Ho & Salimans, 2022) is a simple but effective approach. Intuitively, the classifier-free guidance method tries to let the posterior estimator “know” the label of the data it models. In latent space, it uses one neural network to receive $\mathbf{z}_{t}$ and condition $\mathbf{c}$ to make estimations. We can denote the estimated noise given corresponding condition $\mathbf{c}$ for continuous features as $\bar{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{c})$,
$\bar{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{c})=\hat{\mathbf{\epsilon}}(\mathbf{z}_{t})+w_{g}(\underbrace{\hat{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{c})-\hat{\mathbf{\epsilon}}(\mathbf{z}_{t})}_{\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c})})$ (5)
where $\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c})=\hat{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{c})-\hat{\mathbf{\epsilon}}(\mathbf{z}_{t})$ is guidance of $\mathbf{c}$ and it is scaled by guidance weight $w_{g}$.
Equation 5 can be extended to support multivariate guidance. Especially in tabular synthesis, we want to generate samples conditioned on labels $\mathbf{c}$ and sensitive features $\mathbf{s}$. In this case, we can let our neural network take $\mathbf{z}_{t}$, $\mathbf{c}$, and $\mathbf{s}$ as inputs, resulting in:
$\bar{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{c},\mathbf{s})=\hat{\mathbf{\epsilon}}(\mathbf{z}_{t})+w_{g}(\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c})+\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c},\mathbf{s}))$ (6)
$\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c},\mathbf{s})=\mathbf{\mu}(\mathbf{c},\mathbf{s};w_{s},\lambda)(\hat{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{s})-\hat{\mathbf{\epsilon}}(\mathbf{z}_{t}))$
The sensitive guidance $\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c},\mathbf{s})$ is controlled element-wisely by a “security gate” $\mathbf{\mu}(\mathbf{c},\mathbf{s};w_{s},\lambda)$,
$\mathbf{\mu}(\mathbf{c},\mathbf{s};w_{s},\lambda)=\begin{cases}\phi,&\text{where }\hat{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{c})\ominus\hat{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{s})<\lambda\\ 0,&\text{otherwise }\end{cases}$
$\phi=\max(1,w_{s}|\hat{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{c})-\hat{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{s})|)$
where $w_{s}$ is the sensitive guidance weight. The purpose of $\mathbf{\mu}(\mathbf{c},\mathbf{s};w_{s},\lambda)$ is to keep the alterations caused by the sensitive guidance steady and minimal. It deactivates the sensitive guidance when the difference between the estimation conditioned on $\mathbf{c}$ and the estimation conditioned on $\mathbf{s}$ is larger than a threshold $\lambda$. Furthermore, sensitive guidance is deactivated in the early phase of the diffusion process when $t$ is smaller than a warm-up timestep $\delta$ (a positive integer indicating the timestep at which sensitive guidance starts):
$\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c},\mathbf{s}):=\mathbf{0}\text{ if }t<\delta$
Additionally, a momentum term with momentum weight $w_{m}$ is added to the sensitivity guidance to make guidance direction stable,
$\mathbf{\gamma}_{t}(\mathbf{z}_{t},\mathbf{c},\mathbf{s})=\mathbf{\mu}(\mathbf{c},\mathbf{s};w_{s},\lambda)(\hat{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{s})-\hat{\mathbf{\epsilon}}(\mathbf{z}_{t}))+w_{m}\mathbf{\nu}_{t}$
$\mathbf{\nu}_{t+1}=\beta\mathbf{\nu}_{t}+(1-\beta)\mathbf{\gamma}_{t}$
where $\mathbf{\nu}_{0}=\mathbf{0}$ and $\beta\in[0,1]$. To include not only one sensitive feature, but $N$ sensitive features $\mathbf{S}=\{\mathbf{s}^{(1)},\cdots,\mathbf{s}^{(N)}\}$, we can extend the sensitive guidance as follows:
$\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c},\mathbf{S})=\sum_{i=1}^{N}\mathbf{\mu}(\mathbf{c},\mathbf{s}^{(i)};w_{s},\lambda)(\hat{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{s}^{(i)})-\hat{\mathbf{\epsilon}}(\mathbf{z}_{t}))$
In this way, we can model continuous features given labels $\mathbf{c}$ and a set of sensitive features $\mathbf{S}$. Similarly, the estimated starting point given $\mathbf{c}$ and $\mathbf{S}$ for discrete features can be formulated as follows,
$\bar{\mathbf{x}}_{0}(\mathbf{z}_{t},\mathbf{c},\mathbf{S})=\hat{\mathbf{x}}_{0}(\mathbf{z}_{t})+w_{g}(\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c})+\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c},\mathbf{S}))$ (7)
with $\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c})$ and $\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c},\mathbf{S})$ parameterized by $\hat{\mathbf{x}}_{0}$.
4.2 Backbone
In this research, we model mixed-type tabular data in latent space following a similar setup in (Rombach et al., 2022). The tabular data are encoded into latent space and decoded back with multilayer perceptrons. We choose U-Net (Ronneberger et al., 2015) with transformers as a posterior estimator to predict $\bar{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{c},\mathbf{S})$ and $\bar{\mathbf{x}}_{0}(\mathbf{z}_{t},\mathbf{c},\mathbf{S})$. This choice is based on the hypothesis that the U-Net, augmented by the contextual understanding capabilities of transformers, can effectively manage the heterogeneous nature of tabular data. This integration aims to leverage the U-Net’s ability to capture spatial correlation and the transformers’ strengths in sequence modeling, thereby offering a novel approach to latent space modeling of tabular data.
4.3 Balanced Sampling
After training diffusion models on tabular data paired with labels $\mathbf{c}$ and a set of sensitive features $\mathbf{S}$, we apply multivariate latent guidance for conditional tabular data synthesis. We customize the marginal distributions of the provided labels and sensitive features in the sampling phase. Fair synthetic tabular data are typically anticipated to exhibit the same label distribution as the real data. These data are expected to mirror the label distribution of the real data and exhibit a balanced distribution of sensitive features, which addresses the common issue of label imbalance in real-world datasets. To achieve this, we can compute the empirical distribution of the labels in the actual data and use it together with uniform categorical distributions for each sensitive feature to guide the sampling process. By doing this, we generate balanced synthetic data regarding the sensitive features, and the synthetic data have the same label distribution as the actual data.
5 Experiments
We first introduce experiment datasets, data processing, baselines, and how we assess synthetic tabular data. Then we present computational results on experimental datasets.
5.1 Datasets
We evaluate the effectiveness and fairness of our proposed method using seven tabular datasets for classification tasks. These datasets contain numerical and categorical features, including some sensitive attributes with observed class imbalances. We provide a summary of the datasets, including their corresponding sensitive attributes in Table 1, and offer further details, such as data sources, sensitive class label distributions, etc., in Appendix A. Common sensitive attributes such as sex and race can be found in data sets like UCI Adult, Cardio, Credit Card, Depression, KDD Census, and Law School. Additionally, other sensitive attributes such as marital status and education are present in the Bank Marketing and Credit Card datasets. Each data set is divided into training, validation, and test sets using a 50/25/25 split. Following the training of our diffusion models with multivariate guidance, we proceed to generate synthetic data that adheres to the original label distribution while incorporating a uniform distribution for sensitive features.
Table 1: Experimental datasets
Abbreviation Dataset Size Sensitive Target
AD UCI Adult 48842 Sex, Race Income
BA Bank Marketing 45221 Marital Subscription
CA Cardio 70000 Sex Diagnosis
CR Credit Card 30000 Sex, Education Payment
DE Depression 11000 Sex, Race Diagnosis
KD KDD Census 199523 Sex, Race Income
LA Law School 20800 Sex, Race Admission
5.2 Data Processing and Baselines
Data Processing. Following the data processing steps in (Kotelnikov et al., 2023), we split the numerical and categorical features for each dataset and pre-process them separately. Numerical features are converted through quantile transformation, a non-parametric technique transforming variables into a standard Gaussian distribution based on quantiles, thus normalizing the data and enhancing the robustness to outliers. In preparation for the Markov transition process in the latent diffusion model, we encode categorical features into one-hot vectors and concatenate the resulting vectors with normalized numerical features.
Baselines. In assessing the efficacy of our approach for various datasets, we conduct a comparative analysis against several baselines for tabular data synthesis. Within learning-based methods, we categorize them into diffusion models, generative adversarial networks, and variational autoencoders, selecting the top-performing model from each category with publicly available code. In addition to learning-based techniques, we incorporate SMOTE, a robust and traditional method for tabular data synthesis, into our set of baselines. Specifically, the selected baseline methods for comparison are described in detail as follows.
  • TabDDPM (Kotelnikov et al., 2023): a recent diffusion-based model that has demonstrated to surpass existing methods for synthesizing tabular data. Leveraging multilayer perceptrons as its backbone, TabDDPM excels in learning unbalanced distributions present in the training data.
  • CTGAN (Xu et al., 2019): a tabular data synthesis approach based on generative adversarial networks, providing publicly available code that is accessible and user-friendly in (Patki et al., 2016).
  • TVAE (Xu et al., 2019): a state-of-the-art variational autoencoder specifically designed to generate tabular data with a user-friendly implementation in (Patki et al., 2016).
  • SMOTE (Chawla et al., 2002): a robust interpolation-based method that creates synthetic data points by combining a real data point with its $k$ nearest neighbors from the dataset. In our application, we leverage SMOTE to generate additional tabular data samples, employing interpolation among existing samples with the same label. The implementation of SMOTE is available in (Lemaître et al., 2017).
5.3 Assessing Synthetic Tabular Data
There are two primary approaches to evaluating synthetic tabular data. One approach is a model-based method that involves training machine learning models on synthetic tabular data and evaluating the trained models on validation data extracted from the original dataset. The other approach is a data-based method that operates independently of machine learning models and directly compares synthetic and real tabular data.
Model-Based Evaluation. The computation procedure of model-based evaluation of synthetic tabular data in classification tasks involves the following steps.
  1. 1.
    Train a generative model on the original training set.
  2. 2.
    Generate synthetic data with the generative model.
  3. 3.
    Train classifiers on the synthetic data.
  4. 4.
    Evaluate classifiers on the original validation data.
Specifically, when assessing model accuracy on the original validation data, it is referred to as machine learning efficiency as used in (Choi et al., 2017). Beyond accuracy assessments, the evaluation of synthetic data can extend to fairness scores of classifiers, providing insights into the fairness of the generated data. We evaluate the fairness of machine learning models trained on the synthetic data, focusing on two key aspects: equal allocation and equal performance (Agarwal et al., 2018).
  • Equal allocation, a fundamental principle in model-based fairness evaluations, entails the idea that a model should distribute resources or opportunities proportionally among different groups, irrespective of their affiliation with any privileged group. Equal allocation is quantified using the demographic parity ratio. This ratio reveals the extent of the imbalance by comparing the lowest and highest selection rates (the proportion of examples predicted as positive) between groups.
  • Equal performance is grounded in the principle that a fair model should exhibit consistent performance across all groups. This entails maintaining the same precision level for each group. The equalized odds ratio, which represents the smaller of two ratios comparing true and false positive rates between groups, serves as a metric for assessing performance fairness in this context.
The evaluation of the demographic parity ratio and equalized odds ratio depends on the selection of sensitive features. When evaluating fairness scores in our experiments, we choose the first sensitive feature specified in Table 1. Both the demographic parity ratio and equalized odds ratio are within the range $[0,1]$. The demographic parity ratio highlights the importance of distributing resources equally among different groups, whereas the equalized odds ratio prioritizes impartial decision-making within each specific group. Higher values in either metric indicate progress toward fairness. However, attaining perfect fairness, represented by a score of $1$, may not always be feasible or appropriate depending on the specific context. Our proposed method is expected to have enhanced fairness scores compared to baseline models, and we interpret the results using SHAP feature importance (Lundberg & Lee, 2017).
Data-Based Evaluation. The data-based evaluation directly compares synthetic and real tabular data. To address the privacy concerns when sharing tabular data publicly, the synthetic tabular data is expected not to be a “copy” of the original data. The originality of synthetic data is quantified by utilizing a nearest-neighbor distance metric suggested by (Park et al., 2018), known as the “distance to the closest record” (DCR), calculated as the Euclidean distance between each synthetic sample and its nearest real data point. Larger DCR values indicate that the synthetic data has a greater dissimilarity to the original data. In addition, when evaluating the fairness of synthetic tabular data with a data-based method, we examine the imbalance in class characteristics within sensitive features such as sex or race.
5.4 Computational Results
This section presents the computational results of our evaluation of synthetic tabular data generated with five different random seeds. To ensure the robustness of our model-based evaluations, we employed a diverse set of classic classifiers, including decision trees, random forests, support vector machines, multilayer perceptrons, and XGBoost. We determined the optimal model for each synthetic dataset through meticulous training and evaluation across separate training and validation sets, utilizing five random seeds for each model. The averaged validation accuracy, along with the standard deviation, as shown in Table 2, revealed XGBoost as the dominant performer across most datasets, with random forest excelling in one exceptional case (see fairness scores in Section A.3).
Table 2: The best and average values of validation accuracy on experimental datasets (the averaged validation accuracy $\pm$ the standard deviation)
Dataset Model Choice Best Average
AD XGBoost $0.866_{\pm 0.000}$ $0.795_{\pm 0.041}$
BA XGBoost ${0.900_{\pm 0.000}}$ $0.869_{\pm 0.031}$
CA XGBoost ${0.733_{\pm 0.000}}$ $0.658_{\pm 0.015}$
CR Random Forest ${0.816_{\pm 0.000}}$ $0.757_{\pm 0.018}$
DE XGBoost ${0.941_{\pm 0.000}}$ $0.929_{\pm 0.001}$
KD XGBoost ${0.956_{\pm 0.000}}$ $0.946_{\pm 0.000}$
LA XGBoost ${1.000_{\pm 0.000}}$ $0.942_{\pm 0.004}$
Performance. In our exploration of the machine learning efficiency of synthetic data compared to baselines, we employ two rigorous protocols. The first protocol calculates the average machine learning efficiency across all classifiers, while the second identifies and utilizes the best classifier for each dataset. This multifaceted approach ensures a robust and insightful evaluation. As illustrated in Table 3, our proposed method maintains competitiveness in terms of average machine learning efficiency, even when incorporating a uniformly distributed direction for sensitive features during the sampling phase. It generally outperforms CTGAN and TVAE. Moreover, our method excels against TabDDPM on the majority of datasets. While the interpolation-based method, SMOTE, proves highly competitive and challenging to outperform, our model achieves a machine learning efficiency very close to SMOTE, underscoring its effectiveness. The best machine learning efficiency exhibits a similar pattern with the average machine learning efficiency, with SMOTE and our model emerging as the two most competitive methods.
Table 3: The values of machine learning efficiency averaged on all types of classifiers
Average Accuracy
AD BA CA CR DE KD LA
TabDDPM $0.777_{\pm 0.093}$ $0.874_{\pm 0.017}$ $0.640_{\pm 0.025}$ $0.768_{\pm 0.032}$ $0.920_{\pm 0.001}$ $0.937_{\pm 0.001}$ $0.905_{\pm 0.012}$
CTGAN $0.742_{\pm 0.148}$ $0.873_{\pm 0.019}$ $0.600_{\pm 0.061}$ $0.746_{\pm 0.058}$ $0.867_{\pm 0.050}$ $0.931_{\pm 0.014}$ $0.882_{\pm 0.112}$
TVAE $0.786_{\pm 0.023}$ $0.854_{\pm 0.026}$ $0.613_{\pm 0.058}$ $0.784_{\pm 0.021}$ $0.881_{\pm 0.051}$ $0.938_{\pm 0.008}$ $0.919_{\pm 0.119}$
SMOTE $0.807_{\pm 0.038}$ $0.880_{\pm 0.022}$ $0.632_{\pm 0.074}$ $0.770_{\pm 0.044}$ $0.928_{\pm 0.022}$ $0.939_{\pm 0.013}$ $0.930_{\pm 0.124}$
Ours $0.803_{\pm 0.022}$ $0.880_{\pm 0.011}$ $0.629_{\pm 0.001}$ $0.734_{\pm 0.045}$ $0.926_{\pm 0.004}$ $0.932_{\pm 0.004}$ $0.921_{\pm 0.003}$
Best Accuracy
AD BA CA CR DE KD LA
TabDDPM $0.834_{\pm 0.004}$ $0.895_{\pm 0.003}$ $0.708_{\pm 0.003}$ $0.816_{\pm 0.002}$ $0.937_{\pm 0.003}$ $0.945_{\pm 0.001}$ $0.964_{\pm 0.002}$
CTGAN $0.817_{\pm 0.002}$ ${0.888}_{\pm 0.002}$ ${0.652}_{\pm 0.006}$ ${0.716}_{\pm 0.009}$ ${0.870}_{\pm 0.008}$ ${0.941}_{\pm 0.001}$ ${0.975}_{\pm 0.002}$
TVAE ${0.810}_{\pm 0.004}$ ${0.849}_{\pm 0.008}$ ${0.662}_{\pm 0.006}$ ${0.773}_{\pm 0.006}$ ${0.900}_{\pm 0.003}$ ${0.944}_{\pm 0.001}$ ${0.993}_{\pm 0.001}$
SMOTE $0.853_{\pm 0.000}$ $0.898_{\pm 0.000}$ $0.704_{\pm 0.000}$ $0.817_{\pm 0.001}$ $0.941_{\pm 0.000}$ $0.949_{\pm 0.000}$ $1.000_{\pm 0.000}$
Ours $0.838_{\pm 0.001}$ $0.892_{\pm 0.003}$ $0.709_{\pm 0.004}$ $0.806_{\pm 0.004}$ $0.942_{\pm 0.003}$ $0.930_{\pm 0.003}$ $0.995_{\pm 0.004}$
Privacy. As explained in Section 5.3, we employ DCR values to assess the privacy score of the synthetic tabular data. The median DCR values, obtained across five random seeds, are detailed in Table 4. Remarkably, our method consistently attains a superior privacy score in comparison to both SMOTE and TabDDPM. This suggests that our approach excels in maintaining the distinctiveness and privacy of the synthetic data.
Table 4: The median value of distance to the closest record
AD BA CA CR DE KD LA
TabDDPM $0.179$ $0.195$ $0.021$ $0.123$ $0.520$ $2.089$ $0.176$
CTGAN $0.340$ $0.303$ $0.028$ $0.202$ $0.428$ $2.451$ $0.276$
TVAE $0.253$ $1.415$ $0.032$ $0.208$ $0.443$ $2.018$ $1.414$
SMOTE $0.170$ $0.191$ $0.021$ $0.117$ $0.398$ $2.005$ $0.178$
Ours $0.511$ $0.229$ $0.023$ $0.255$ $0.430$ $2.831$ $0.216$
Fairness. A key contribution of our work is the generation of balanced tabular data while considering sensitive features. To substantiate our claim and provide empirical evidence, we conduct a thorough assessment of the fairness of our proposed model. After choosing the best-performing classifier for each dataset, we assess its fairness by evaluating two metrics: demographic parity ratio and equalized odds ratio, and the results are presented in Table 5. Our approach outperforms baseline methods notably in terms of equal allocation, as evidenced by higher values of demographic parity ratio across nearly all datasets. For instance, on the KDD Census dataset, our method achieves a demographic parity ratio of $0.613$, surpassing baseline methods like SMOTE and TabDDPM by around fifty percent, while SMOTE only attains $0.141$, and TabDDPM reaches $0.097$. Similarly, on the UCI Adult dataset, our model demonstrates a demographic parity ratio of $0.529$, surpassing the ratios of $0.306$ by SMOTE and $0.312$ by TabDDPM. In addition, our model demonstrates competitiveness in terms of equal performance metrics, as indicated by the equalized odds ratio. Particularly noteworthy are datasets such as UCI Adult and KDD Census, where our model achieves an equalized odds ratio of $0.641$ and $0.884$, respectively, surpassing baseline models that register less than $0.250$ in these metrics. Furthermore, employing equitable synthetic data to train classifiers can yield higher fairness scores compared to training with the original data, as illustrated in Table 7. In contrast, baseline models generally achieve comparable fairness scores.
Table 5: The values of demographic parity ratio and equalized odds ratio with the best-performing classifier
Demographic Parity Ratio
AD BA CA CR DE KD LA
TabDDPM $0.312_{\pm 0.033}$ $0.590_{\pm 0.070}$ $0.957_{\pm 0.036}$ $0.791_{\pm 0.063}$ $0.634_{\pm 0.020}$ $0.097_{\pm 0.032}$ $0.858_{\pm 0.007}$
TabGAN ${0.188}_{\pm 0.039}$ ${0.749}_{\pm 0.060}$ ${0.760}_{\pm 0.027}$ ${0.915}_{\pm 0.051}$ ${0.706}_{\pm 0.021}$ ${0.112}_{\pm 0.037}$ ${0.844}_{\pm 0.006}$
TabVAE ${0.249}_{\pm 0.017}$ ${0.624}_{\pm 0.045}$ ${0.965}_{\pm 0.011}$ ${0.740}_{\pm 0.088}$ ${0.664}_{\pm 0.012}$ ${0.092}_{\pm 0.027}$ ${0.847}_{\pm 0.006}$
SMOTE ${0.306}_{\pm 0.000}$ ${0.625}_{\pm 0.000}$ ${0.975}_{\pm 0.000}$ ${0.787}_{\pm 0.000}$ ${0.671}_{\pm 0.000}$ ${0.141}_{\pm 0.000}$ ${0.851}_{\pm 0.000}$
Ours $0.529_{\pm 0.046}$ $0.839_{\pm 0.098}$ $0.963_{\pm 0.017}$ $0.828_{\pm 0.048}$ $0.706_{\pm 0.005}$ $0.613_{\pm 0.057}$ $0.862_{\pm 0.007}$
Equalized Odds Ratio
AD BA CA CR DE KD LA
TabDDPM $0.247_{\pm.0.033}$ $0.496_{\pm.0.077}$ $0.919_{\pm.0.068}$ $0.680_{\pm.0.081}$ $0.764_{\pm.0.110}$ $0.134_{\pm.0.050}$ $0.593_{\pm.0.484}$
CTGAN ${0.125}_{\pm 0.047}$ ${0.760}_{\pm 0.074}$ ${0.669}_{\pm 0.038}$ ${0.895}_{\pm 0.048}$ ${0.815}_{\pm 0.062}$ ${0.178}_{\pm 0.067}$ ${0.611}_{\pm 0.172}$
TVAE ${0.205}_{\pm 0.023}$ ${0.621}_{\pm 0.046}$ $0.950_{\pm 0.027}$ ${0.667}_{\pm 0.115}$ ${0.825}_{\pm 0.067}$ ${0.126}_{\pm 0.046}$ ${0.562}_{\pm 0.175}$
SMOTE ${0.231}_{\pm 0.000}$ ${0.593}_{\pm 0.000}$ ${0.939}_{\pm 0.000}$ ${0.683}_{\pm 0.000}$ ${0.827}_{\pm 0.000}$ ${0.202}_{\pm 0.000}$ ${1.000}_{\pm 0.000}$
Ours $0.641_{\pm 0.108}$ $0.760_{\pm 0.043}$ $0.901_{\pm 0.035}$ $0.802_{\pm 0.056}$ $0.254_{\pm 0.089}$ $0.884_{\pm 0.055}$ $0.170_{\pm 0.340}$
Figure 2: Comparison in SHAP values on KDD dataset: balanced scenario (top) vs. unbalanced scenario (bottom) using XGBoost
In Figure 2, we illustrate the feature importance through absolute SHAP values when making predictions with XGBoost on the KDD Census dataset. On the top, we present the scenario where synthetic data generated by our approach exhibits a uniformly distributed sensitive feature, sex, across all samples. In contrast, the figure on the bottom showcases data generated by TabDDPM with biased sampling in the sex feature. It is evident that in the data generated by TabDDPM, the variable sex significantly influences decision-making, while our method mitigates the impact of sex on the final decision to a considerable extent.
Apart from fairness evaluation dependent on machine learning models, we also analyze the class distribution within the sensitive attributes. For example, in the UCI Adult dataset, the distribution of sensitive attributes has an inherent bias. In contrast, the synthetic data generated through our approach exhibits more balanced distributions, as depicted in Figure 3, where the y-axis represents the percentage of the respective attribute. We present visualizations of class imbalance in sensitive attributes for the rest of experimental datasets in Figure 4.
Figure 3: Comparison in the real vs. synthetic distribution of sensitive attributes in the UCI Adult dataset
5.5 Limitations and Discussion
Our proposed method is designed to generate balanced synthetic tabular data considering sensitive attributes. Empirically, classifiers trained on such balanced synthetic tabular data demonstrate higher fairness scores at equal allocation metrics indicated by demographic parity ratios. However, there is no guarantee that equal performance, reflected by equalized odds ratios, will always improve. We use U-Net with attention layers as the backbone neural network for the posterior estimator in the diffusion framework, but it is time-consuming. Despite our model exhibiting superior machine learning efficiency and higher fairness scores compared to other deep generative models in our experiments, the computational time is approximately five times longer. In the future, we may explore the incorporation of constraints to ensure equal performance and seek ways to reduce computational costs.
6 Conclusion
In this work, we propose a novel diffusion model framework for mixed-type tabular data conditioned on both outcome and sensitive feature variables. Our approach leverages a multivariate guidance mechanism and performs balanced sampling considering sensitive features while ensuring a fair representation of the generated data. Extensive experiments on real-world datasets containing sensitive demographics demonstrate that our model achieves competitive performance and superior fairness compared to existing baselines.
Impact Statements
The objective of this study is to make progress in the area of fair machine learning. Tabular datasets sometimes contain inherent bias, such as imbalanced distributions in sensitive attributes. Training machine learning models on biased datasets may result in decisions that could negatively affect minority groups. By mitigating biases present in tabular datasets through the generation of equitable synthetic data, our approach contributes to fostering equitable decision-making processes across industries such as finance, healthcare, and employment. Moreover, in instances where sharing datasets becomes necessary, the utilization of fair synthetic data ensures the preservation of user privacy and avoids causing emotional distress to minority groups. An adverse possibility is that bad individuals could manipulate distributions in sensitive attributes to generate biased (even stronger than original) data to harm minority groups.
Appendix A Details about Datasets
A.1 Data Sources
A.1.1 UCI Adult
The UCI Adult dataset (Kohavi et al., 1996) is widely used as a benchmark for exploring fairness and bias in machine learning. It contains 48842 data points. Each data point has 14 attributes and a binary target variable indicating whether an individual earns over fifty thousand dollars annually. The dataset encompasses employment, education, and demographic information, and sensitive features are sex and race. Download it from OpenML.
A.1.2 Bank Marketing
The Bank Marketing dataset is collected from direct marketing campaigns of a Portuguese banking institution. Comprising 45211 instances with 16 features, the predicted outcome is a binary variable indicating whether a client subscribed to a term deposit. Demographic, economic, and past marketing campaign data are included in features, and marital status is sensitive. Download it from OpenML.
A.1.3 Cardio
The Cardio dataset focuses on diagnosing cardiovascular disease (yes or no) and it contains 70000 instances. It incorporates 11 features, which contain demographics and factual information (blood pressure, cholesterol levels, etc.), as well as self-reported labels such as whether smoking or not. The sensitive feature is sex. Download it from OpenML.
A.1.4 Credit Card
The Credit Card dataset (Yeh & Lien, 2009) explores customer default payments (yes or no) among credit card clients in Taiwan, and it contains 30000 data points. Each data point has 23 features including the information of demographics and financial records (monthly payments, etc.). The sensitive features are sex and education. Download it from OpenML.
A.1.5 Depression
The Depression dataset contains demographic and clinical data (concentration, rumination, etc.) to predict a diagnosis of depression. It has 11000 instances, each with 18 features, and sensitive features are sex and race. Download it from OpenML.
A.1.6 KDD Census
The KDD Census dataset is an expanded version of the UCI Adult dataset. It involves 199523 instances, each featuring 41 attributes and a binary target variable indicating income over fifty thousand dollars per year. The dataset covers employment, education, and demographic details, with sex and race as sensitive features. Download it from OpenML.
A.1.7 Law School
The Law School dataset presents a survey on the law school admission process. It has 20800 instances. Each instance has 12 attributes and a binary target variable indicating admission to law school. Demographic and academic information, such as GPA, are included in features. The sensitive features are sex and race. Download it from OpenML.
A.2 Number of Features
Dataset Number of Numerical Features Number of Categorical Features
UCI Adult 6 8
Bank Marketing 7 9
Cardio 5 6
Credit Card 14 9
Depression 14 4
KDD Census 7 34
Law School 6 6
Table 6: The number of numerical attributes and categorical attributes in experimental datasets
A.3 Evaluation of Best-Performing Classifiers on Original Datasets
Dataset Model Choice Accuracy Demographic Parity Ratio Equalized Odds Ratio
UCI Adult XGBoost $0.866$ $0.353$ $0.273$
Bank Marketing XGBoost $0.900$ $0.619$ $0.540$
Cardio XGBoost $0.733$ $0.992$ $0.989$
Credit Card Random Forest $0.816$ $0.778$ $0.834$
Depression XGBoost $0.941$ $0.696$ $0.644$
KDD Census XGBoost $0.956$ $0.151$ $0.170$
Law School XGBoost $1.000$ $0.825$ $1.000$
Table 7: The values of accuracy and fairness scores of the best-performing classifier on experimental datasets
A.4 Class Imbalance
(a) Bank Marketing
(b) Cardio
(c) Credit Card
(d) Depression
(e) KDD Census
(f) Law School
Figure 4: Visualizations of class imbalance in sensitive attributes in experimental datasets except UCI Adult

### v3

@@
 Abstract
-Diffusion models have emerged as a robust framework for various generative tasks, such as image and audio synthesis, and have also demonstrated a remarkable ability to generate mixed-type tabular data comprising both continuous and discrete variables.
-However, current approaches to training diffusion models on mixed-type tabular data tend to inherit the imbalanced distributions of features present in the training dataset, which can result in biased sampling.
-In this research, we introduce a fair diffusion model designed to generate balanced data on sensitive attributes.
-We present empirical evidence demonstrating that our method effectively mitigates the class imbalance in training data while maintaining the quality of the generated samples.
-Furthermore, we provide evidence that our approach outperforms existing methods for synthesizing tabular data in terms of performance and fairness.
-Machine Learning, Diffusion Models
+Diffusion models have emerged as a robust framework for various generative tasks, including tabular data synthesis.
+However, current tabular diffusion models tend to inherit bias in the training dataset and generate biased synthetic data, which may influence discriminatory actions.
+In this research, we introduce a novel tabular diffusion model that incorporates sensitive guidance to generate fair synthetic data with balanced joint distributions of the target label and sensitive attributes, such as sex and race.
+The empirical results demonstrate that our method effectively mitigates bias in training data while maintaining the quality of the generated samples.
+Furthermore, we provide evidence that our approach outperforms existing methods for synthesizing tabular data on fairness metrics such as demographic parity ratio and equalized odds ratio, achieving improvements of over $10\%$.
+Our implementation is available at https://github.com/comp-well-org/fair-tab-diffusion.
 1 Introduction
-Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) are a class of energy-based generative models utilizing a stochastic process to transform desired data samples from noise iteratively and have demonstrated remarkable performance in generative modeling for various tasks, especially text-to-image synthesis.
+Tabular data synthesis, particularly when involving mixed-type variables, presents unique challenges due to the need to simultaneously handle continuous and discrete features.
+Recent advances in diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) are now being leveraged to tackle these challenges.
+Diffusion models have demonstrated remarkable performance in generative modeling for various tasks, especially text-to-image synthesis.
 Researchers use diffusion frameworks to model images paired with text descriptions and show promising results (Nichol et al., 2021; Saharia et al., 2022).
 Advancing further from diffusion models in pixel space, latent diffusion frameworks optimize image representation in a low-dimensional space, thereby enhancing synthetic image quality and efficiency (Rombach et al., 2022; Ramesh et al., 2022).
-The impressive results of diffusion models in image synthesis extended to the generation of sequential data such as audio and time series (Kong et al., 2020; Rasul et al., 2021; Li et al., 2022).
-This advancement has gradually overshadowed generative adversarial networks (previous research spot in generative modeling), as diffusion models consistently demonstrate superior performance across diverse tasks (Dhariwal & Nichol, 2021; Stypułkowski et al., 2023; Mazé & Ahmed, 2023).
+The impressive results of diffusion models in image synthesis extended to generating sequential data such as audio and time series (Kong et al., 2020; Rasul et al., 2021; Li et al., 2022).
+This advancement has gradually overshadowed generative adversarial networks, which were previously a significant focus in generative modeling, as diffusion models consistently demonstrate superior performance across diverse tasks (Dhariwal & Nichol, 2021; Stypułkowski et al., 2023; Mazé & Ahmed, 2023).
 Researchers have recently been motivated by the versatility of diffusion models to explore the potential in synthesizing tabular data (Kotelnikov et al., 2023; He et al., 2023).
 Tabular data often involves sensitive user information, which raises significant privacy concerns when used for model training.
-Moreover, such datasets suffer from limitations in size and imbalanced class distribution (Jesus et al., 2022), which leads to challenges in training robust and fair machine learning models.
+Moreover, such datasets suffer from limitations in size and imbalanced class distribution (Jesus et al., 2022), leading to challenges in training robust and fair machine learning models.
 To address these issues, there is a growing demand for effective tabular data synthesis.
 However, generative modeling of tabular data is challenging due to the combination of discrete and continuous features and their varying value distributions.
-Interpolation-based techniques, such as SMOTE (Chawla et al., 2002), are simple but effective and challenging to beat.
+Interpolation-based techniques, such as SMOTE (Chawla et al., 2002), are simple and effective but they pose privacy concerns as they generate new data points through direct interpolation of existing data.
 Researchers have used adversarial generative networks and variational autoencoders to generate tabular data with promising results (Ma et al., 2020; Xu et al., 2019).
 Research on diffusion models for tabular data is ongoing, and the primary challenge is to handle continuous and discrete features simultaneously.
-In (Austin et al., 2021), the authors suggested using diffusion frameworks for discrete features with Markov transition matrices.
-Recent developments such as TabDDPM (Kotelnikov et al., 2023), which combines Markov transition functions (called diffusion kernels) for continuous and discrete features, mark significant progress in this field and underscore the potential of diffusion models as a comprehensive solution for tabular data synthesis.
+In Austin et al.
+(2021), the authors suggest using diffusion frameworks for discrete features with Markov transition matrices.
+Recent developments such as TabDDPM (Kotelnikov et al., 2023) and CoDi (Lee et al., 2023), which combine Markov transition functions (called diffusion kernels) for continuous and discrete features, mark significant progress in this field and underscore the potential of diffusion models as a comprehensive solution for tabular data synthesis.
 Despite the success of diffusion models in various generative tasks, ethical concerns regarding potential bias and unsafe content in the synthetic data persist.
 Generative models are inherently data-driven, thereby propagating these biases into the synthetic data.
-Studies such as (Friedrich et al., 2023; Schramowski et al., 2023) have shown that latent diffusion models, particularly in text-to-image synthesis, can create images biased toward certain privileged groups and may contain inappropriate elements, such as nudity and violence.
+Studies such as fair diffusion (Friedrich et al., 2023) and safe latent diffusion (Schramowski et al., 2023) have shown that latent diffusion models, particularly in text-to-image synthesis, can create images biased toward certain privileged groups and may contain inappropriate elements, such as nudity and violence.
 In response, they propose methods to apply hazardous or sensitive guidance in latent space and then eliminate bias and offensive elements in the synthetic images.
@@
 For instance, some datasets exhibit significant imbalances in demographic groups regarding sensitive attributes such as sex and race (Le Quy et al., 2022).
-However, in our review of the literature, research on the safety and fairness of diffusion models in areas beyond text-to-image synthesis is limited.
-Although diffusion models have shown remarkable performance in tabular data synthesis (Sattarov et al., 2023; He et al., 2023; Kotelnikov et al., 2023), they are prone to learning and replicating these biases in training data.
-To address this issue, we propose a novel diffusion framework to accurately model distributions of mixed-type tabular data, with a specific focus on outcomes and sensitive attributes.
-We employ balanced sampling techniques to reduce disparities in group representation within the synthetic data.
+Training machine learning models on biased tabular data can result in discriminatory outcomes for underrepresented groups.
+However, in our literature review, research on the safety and fairness of diffusion models in areas beyond text-to-image synthesis is limited.
+Although diffusion models have shown remarkable performance in tabular data synthesis (Sattarov et al., 2023; He et al., 2023; Kotelnikov et al., 2023; Lee et al., 2023; Kim et al., 2022), they are prone to learning and replicating these biases in training data.
+Existing tabular diffusion models are typically either unconditional or conditioned on a single target variable.
+For example, CoDi (Lee et al., 2023) leverages two interconnected diffusion models to separately process continuous and discrete variables while maintaining their inter-variable correlations through co-evolutionary conditioning and contrastive learning.
+Although CoDi efficiently captures inter-variable relationships, its conditioning mechanism remains restricted to single-label guidance.
+As a result, CoDi does not explicitly address fairness considerations in the presence of sensitive attributes.
+In contrast, we introduce a novel tabular diffusion model that approximates mixed-type tabular data distributions, conditioned on both outcomes and multiple sensitive attributes.
+Our method extends tabular diffusion models from label-only conditioning to multivariate feature-level conditioning for enhancing fairness.
+To promote fairness, our approach generates data with balanced joint distributions of the target label and sensitive attributes.
+Specifically, our approach integrates sensitive guidance into the tabular diffusion model and extends it to manage multiple sensitive attributes, such as sex and race.
+We utilize a U-Net architecture to approximate the data distributions in the latent space during the reverse diffusion process conditioned on various variables, including the target label and sensitive attributes.
+To ensure fairness while preserving the integrity of the original data, we apply the element-wise comparison technique from safe latent diffusion to regulate the sensitive guidance within an appropriate range.
+Additionally, we employ balanced sampling techniques to reduce disparities in group representation within the synthetic data.
+Given that our method leverages diffusion frameworks tailored for tabular data and supports multiple variables for conditional generation of fair tabular data, it also has the potential for extension to multi-modal data generation with fairness considerations using diffusion models.
+In summary, our main contributions are as follows: We employ balanced sampling techniques to reduce disparities in group representation within the synthetic data.
 In summary, our main contributions are as follows:
   1. 1.
-    We introduce a diffusion-based framework tailored to learn distributions of mixed-type tabular data given multiple attributes.
+    We introduce, to our knowledge, the first diffusion-based framework tailored to learn distributions of mixed-type tabular data with the extension from label-only conditioning to multivariate feature-level conditioning for fairness.
   2. 2.
@@
   3. 3.
-    Our model surpasses existing models in specific areas of performance and fairness, as evidenced through comprehensive evaluations.
-The remainder of this paper is organized as follows: Section 2 provides an overview of fairness-aware machine learning and diffusion models.
-Section 3 provides an overview of mixed-type modeling with diffusion models.
-In Section 4, we detail our approach to multivariate conditioning and balanced sampling.
+    Our model demonstrates strong performance in terms of fidelity and diversity of synthetic data, while surpassing existing models in fairness metrics such as demographic parity ratio and equalized oddes ratio by more than $10\%$ on average across three experimental datasets.
+The remainder of this paper is organized as follows: Section 2 provides an overview of related work on fairness-aware machine learning and diffusion models.
+Section 3 discusses the relevant background of mixed-type modeling with diffusion models.
+In Section 4, we present our approach to multivariate conditioning and balanced sampling.
 In Section 5, we illustrate the experiments and present the results.
+Section 6 discusses the limitations of our method.
 Finally, in Section 7, we conclude our findings and discuss their implications.
 2 Related Work
+2.1 Bias in Machine Learning
 The concept of bias in machine learning that involves assigning significance to specific features to enhance overall generalization is essential for the effective performance of models (Mehrabi et al., 2021).
@@
 Lastly, post-processing methods (Hardt et al., 2016; Pleiss et al., 2017), implemented after training machine learning models on collected datasets, directly override the predicted labels to improve fairness.
-Fairness-aware generative models can be classified as pre-processing methods.
-In one of such methods proposed to generate fair data, (Xu et al., 2018) introduced fairness-aware adversarial generative networks that employ a fair discriminator to maintain equality in the joint probability of features and labels conditioned on subgroups within sensitive attributes.
-However, this approach does not effectively tackle the imbalance in sensitive classes within synthetic data, and diffusion-based tabular data synthesis methods such as TabDDPM (Kotelnikov et al., 2023) excel in producing synthetic data of superior quality.
+2.2 Fair Data Synthesis
+Fairness-aware generative models typically operate as pre-processing methods.
+One such method proposed by Xu et al.
+(2018) introduced fairness-aware adversarial generative networks that employ a fair discriminator to maintain equality in the joint probability of features and labels conditioned on subgroups within sensitive attributes.
+Similarly, DECAF (Van Breugel et al., 2021), a causally-aware GAN-based framework, incorporates structural causal models into the generator to ensure fairness, enabling inference-time debiasing and compatibility with multiple fairness definitions.
+However, these approaches do not effectively tackle the imbalance in sensitive classes within synthetic data, and diffusion-based tabular data synthesis methods such as Kotelnikov et al.
+(2023); Lee et al.
+(2023); Zhang et al.
+(2023) excel in producing synthetic data of superior quality.
+Fair Class Balancing (Yan et al., 2020), an oversampling algorithm based on SMOTE, demonstrates impressive results but lacks flexibility in manipulating data distributions.
+As an interpolation-based method, it struggles with data diversity and is not well-suited for multi-modal generation.
+Recently, Fair4Free (Sikder et al., 2024) introduced a novel approach to fairness-aware data synthesis using data-free distillation, which trains a smaller student model to generate fair synthetic data while preserving privacy and outperforming state-of-the-art methods in fairness, utility, and data quality.
+2.3 Fair/Safe Diffusion Models
 With impressive capabilities in generative modeling, diffusion models are a class of deep generative models that utilize a Markov process that starts from noise and ends with the target data distribution.
-Although diffusion models are widely studied, the research on fairness and safety of diffusion models is limited, and existing works (Schramowski et al., 2023; Friedrich et al., 2023) mainly focus on text-to-image synthesis tasks.
-Researchers in (Friedrich et al., 2023) utilize sensitive content, such as sex and race, within text prompts to guide training diffusion models.
+Diffusion models have demonstrated superior performance over large language models in modeling structured data like images and tables, as they are specifically designed to handle mixed-type features and maintain complex inter-dependencies (Zhang et al., 2023).
+Although diffusion models are widely studied, research on the fairness and safety of diffusion models is limited, and existing works (Schramowski et al., 2023; Friedrich et al., 2023) mainly focus on text-to-image synthesis tasks.
+Friedrich et al.
+(2023) utilize sensitive content, such as sex and race, within text prompts to guide training diffusion models.
 Subsequently, during the sampling phase, a uniform distribution of sensitive attributes is applied to generate images that are fair and do not exhibit preference towards privileged groups.
 Similarly, (Schramowski et al., 2023) uses unsafe text, such as nudity and violence, to guide diffusion models and remove unsafe content in the sampling phase.
-In addition, this method applies some tricks to remove unsafe content from synthetic images while keeping changes minimal.
+In addition, this method applies some techniques to remove unsafe content from synthetic images while keeping changes minimal.
 However, these fairness-aware diffusion models have not been adapted to tabular data synthesis, even though bias commonly exists in tabular datasets (Le Quy et al., 2022).
@@
 Our method can generate balanced tabular data considering multiple sensitive attributes and subsequently achieve improved fairness scores without compromising the quality of the synthetic data.
-3 Background
+3 Diffusion Models
 Diffusion models, taking inspiration from thermodynamics, are probabilistic generative models that operate under the Markov assumption.
@@
 $p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1}|\hat{\mathbf{\mu}}(\mathbf{x}_{t}),\hat{\mathbf{\Sigma}}(\mathbf{x}_{t}))$
-Researchers in (Ho et al., 2020) set the estimated covariance matrix $\hat{\mathbf{\Sigma}}(\mathbf{x}_{t})$ as diagonal matrix with constant values $\beta_{t}$ or $\tilde{\beta}_{t}$, and calculate the estimated mean $\hat{\mathbf{\mu}}(\mathbf{x}_{t})$ as follows,
+Ho et al.
+(2020) set the estimated covariance matrix $\hat{\mathbf{\Sigma}}(\mathbf{x}_{t})$ as a diagonal matrix with constant values $\beta_{t}$ or $\tilde{\beta}_{t}$, and calculate the estimated mean $\hat{\mathbf{\mu}}(\mathbf{x}_{t})$ as follows,
 $\hat{\mathbf{\mu}}(\mathbf{x}_{t})=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\mathbf{\epsilon}}(\mathbf{x}_{t}))$ (2)
@@
 $\tilde{\mathbf{\theta}}=[\alpha_{t}\mathbf{x}_{t}+(1-\alpha_{t})/K]\odot[\bar{\alpha}_{t-1}\mathbf{x}_{0}+(1-\bar{\alpha}_{t-1})/K]$
-The posterior estimator for discrete features is preferably parameterized by the estimated starting point $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$ as suggested by (Hoogeboom et al., 2021).
+The posterior estimator for discrete features is preferably parameterized by the estimated starting point $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$ as suggested by Hoogeboom et al. (2021).
 $p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))$ (3)
@@
 The estimated original input $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$ in Equation 3 can be reformulated as $\hat{\mathbf{x}}_{0}(\mathbf{z}_{t})$.
+3.4 Classifier-Free Guidance
+Classifier-free guidance (Ho & Salimans, 2022) is a mechanism used in diffusion models to condition the generation process on auxiliary information without requiring an external classifier.
+Below, we provide a mathematical overview of this method.
+Let $x_{t}$ denote the noisy data at timestep $t$ in the reverse diffusion process, and let $c$ represent the conditional information (e.g., a class label).
+The reverse diffusion process aims to estimate the posterior $p(x_{t-1}|x_{t},c)$, which is parameterized by a neural network $\epsilon_{\theta}$.
+Classifier-free guidance interpolates between conditional and unconditional estimates of the noise, as shown below,
+$\epsilon(x_{t},c)=\epsilon_{\theta}(x_{t})+w_{g}(\epsilon_{\theta}(x_{t},c)-\epsilon_{\theta}(x_{t}))$
+where $\epsilon_{\theta}(x_{t})$ is the unconditional noise estimate, $\epsilon_{\theta}(x_{t},c)$ is the conditional noise estimate, and $w_{g}$ is the guidance weight, controlling the influence of the conditional term.
+The term $\epsilon_{\theta}(x_{t},c)-\epsilon_{\theta}(x_{t})$ represents the guidance signal derived from the condition $c$.
+By scaling this signal with $w_{g}$, the model adjusts the strength of conditioning during generation.
+Larger values of $w_{g}$ enforce stronger adherence to the condition, while smaller values reduce its influence, yielding outputs closer to the unconditional distribution.
+For tasks involving multiple conditions, the guidance term can be extended to support multivariate conditioning.
+Suppose we have $N$ conditions $c_{1},c_{2},\ldots,c_{N}$, the guided noise estimate can be formulated as,
+$\epsilon(x_{t},c_{1},c_{2},\ldots,c_{N})=\epsilon_{\theta}(x_{t})+\sum_{i=1}^{N}w_{g_{i}}\gamma(x_{t},c_{i})$
+where $\gamma(x_{t},c_{i})=\epsilon_{\theta}(x_{t},c_{i})-\epsilon_{\theta}(x_{t})$ represents the guidance signal for the $i$-th condition, and $w_{g_{i}}$ represents the $i$-th guidance weight.
+The overall guidance is computed as the sum of these individual contributions across all conditions.
 4 Methods
@@
 The estimated data point is $\hat{\mathbf{x}}$.
+Our model can generate fair synthetic data by leveraging sensitive guidance to ensure a balanced joint distribution of the target label and sensitive attributes.
 4.1 Multivariate Latent Guidance
-Conditional data generation is important in many tasks like text-to-image synthesis.
 To achieve conditional data generation with diffusion models, classifier-free guidance (Ho & Salimans, 2022) is a simple but effective approach.
@@
 $\bar{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{c},\mathbf{s})=\hat{\mathbf{\epsilon}}(\mathbf{z}_{t})+w_{g}(\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c})+\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c},\mathbf{s}))$ (6)
+Using the mechanism of multivariate classifier-free guidance introduced in Section 3.4, Equation 6 extends label-only conditioning in classifier-free guidance with multivariate feature-level conditioning.
+The term $\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c})$ is the guidance term for the target label, representing the adjustment to the noise estimate based on the label condition.
+The term $\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c},\mathbf{s})$ is the sensitive guidance term, incorporating adjustments based on the sensitive attribute $\mathbf{s}$, and the label condition $\mathbf{c}$ here is used as a reference to restrict the influence of the sensitive guidance, ensuring fairness without distorting the primary label distribution.
+Since the primary conditional guidance is the target label, incorporating sensitive guidance from multiple attributes should influence the synthetic data without significantly altering it compared to using only the target label as guidance.
+To achieve this, the sensitive guidance $\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c},\mathbf{s})$ is controlled element-wisely by a “security gate” $\mathbf{\mu}(\mathbf{c},\mathbf{s};w_{s},\lambda)$,
 $\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c},\mathbf{s})=\mathbf{\mu}(\mathbf{c},\mathbf{s};w_{s},\lambda)(\hat{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{s})-\hat{\mathbf{\epsilon}}(\mathbf{z}_{t}))$
-The sensitive guidance $\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c},\mathbf{s})$ is controlled element-wisely by a “security gate” $\mathbf{\mu}(\mathbf{c},\mathbf{s};w_{s},\lambda)$,
 $\mathbf{\mu}(\mathbf{c},\mathbf{s};w_{s},\lambda)=\begin{cases}\phi,&\text{where }\hat{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{c})\ominus\hat{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{s})<\lambda\\ 0,&\text{otherwise }\end{cases}$
 $\phi=\max(1,w_{s}|\hat{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{c})-\hat{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{s})|)$
-where $w_{s}$ is the sensitive guidance weight.
+where $\ominus$ is element-wise subtraction, and $w_{s}$ is the sensitive guidance weight.
 The purpose of $\mathbf{\mu}(\mathbf{c},\mathbf{s};w_{s},\lambda)$ is to keep the alterations caused by the sensitive guidance steady and minimal.
@@
 $\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c},\mathbf{S})=\sum_{i=1}^{N}\mathbf{\mu}(\mathbf{c},\mathbf{s}^{(i)};w_{s},\lambda)(\hat{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{s}^{(i)})-\hat{\mathbf{\epsilon}}(\mathbf{z}_{t}))$
-In this way, we can model continuous features given labels $\mathbf{c}$ and a set of sensitive features $\mathbf{S}$.
+By this approach, we can model continuous features given labels $\mathbf{c}$ and a set of sensitive features $\mathbf{S}$.
 Similarly, the estimated starting point given $\mathbf{c}$ and $\mathbf{S}$ for discrete features can be formulated as follows,
@@
 This choice is based on the hypothesis that the U-Net, augmented by the contextual understanding capabilities of transformers, can effectively manage the heterogeneous nature of tabular data.
-This integration aims to leverage the U-Net’s ability to capture spatial correlation and the transformers’ strengths in sequence modeling, thereby offering a novel approach to latent space modeling of tabular data.
+This integration leverages the U-Net’s ability to capture spatial correlation and the transformers’ strengths in sequence modeling, thereby offering a novel approach to latent space modeling of tabular data.
 4.3 Balanced Sampling
-After training diffusion models on tabular data paired with labels $\mathbf{c}$ and a set of sensitive features $\mathbf{S}$, we apply multivariate latent guidance for conditional tabular data synthesis.
-We customize the marginal distributions of the provided labels and sensitive features in the sampling phase.
-Fair synthetic tabular data are typically anticipated to exhibit the same label distribution as the real data.
-These data are expected to mirror the label distribution of the real data and exhibit a balanced distribution of sensitive features, which addresses the common issue of label imbalance in real-world datasets.
-To achieve this, we can compute the empirical distribution of the labels in the actual data and use it together with uniform categorical distributions for each sensitive feature to guide the sampling process.
-By doing this, we generate balanced synthetic data regarding the sensitive features, and the synthetic data have the same label distribution as the actual data.
+After training diffusion models to approximate the distribution of the training data conditioned on labels $\mathbf{c}$ and sensitive features $\mathbf{S}$, we can generate samples from this conditional distribution and align them with any desired target distribution by adjusting the conditioning variables.
+Specifically, during sampling, the denoising neural network utilizes guidance from the target label and sensitive attributes.
+The joint distribution of the target label and sensitive attributes can be customized.
+In our bias mitigation framework, we expect the synthetic data to preserve the same label distribution as the real data, while making the sensitive attributes nearly independent of the target label.
+To achieve this, we first compute the empirical label distribution from the real data and apply uniform categorical distributions for each sensitive feature.
+This results in a balanced joint distribution of the target label and sensitive features, which serves as input for the conditional generation in our diffusion model.
+Then our diffusion model iteratively denoises the Gaussian noise with label and sensitive guidance to generate fair tabular data.
 5 Experiments
-We first introduce experiment datasets, data processing, baselines, and how we assess synthetic tabular data. Then we present computational results on experimental datasets.
+In this section, we first introduce experimental datasets, data processing, baselines, and how we assess synthetic tabular data.
+Then we present computational results on experimental datasets.
+The training settings are presented in Appendix B.
 5.1 Datasets
-We evaluate the effectiveness and fairness of our proposed method using seven tabular datasets for classification tasks.
-These datasets contain numerical and categorical features, including some sensitive attributes with observed class imbalances.
-We provide a summary of the datasets, including their corresponding sensitive attributes in Table 1, and offer further details, such as data sources, sensitive class label distributions, etc., in Appendix A.
-Common sensitive attributes such as sex and race can be found in data sets like UCI Adult, Cardio, Credit Card, Depression, KDD Census, and Law School.
-Additionally, other sensitive attributes such as marital status and education are present in the Bank Marketing and Credit Card datasets.
-Each data set is divided into training, validation, and test sets using a 50/25/25 split.
+We evaluate the effectiveness and fairness of our proposed method using three binary classification tabular datasets: Adult (Kohavi et al., 1996), Bank Marketing (Moro et al., 2014), and COMPAS (Larson et al., 2016).
+These datasets contain numerical and categorical features, including some sensitive attributes with observed class imbalances or strong correlations to the target variable.
+We provide a summary of the datasets, including their corresponding sensitive attributes in Table 1, and offer further details in Appendix A.
+Common sensitive attributes such as sex and race can be found in Adult and COMPAS.
+As mentioned in (Rajabi & Garibay, 2022), the age group in the Bank Marketing dataset is considered sensitive, with individuals below 25 classified as “young” and those above 25 classified as “old”.
+In our experiments, each data set is divided into training, validation, and test sets using a 50/25/25 split.
 Following the training of our diffusion models with multivariate guidance, we proceed to generate synthetic data that adheres to the original label distribution while incorporating a uniform distribution for sensitive features.
-Table 1: Experimental datasets
-Abbreviation Dataset Size Sensitive Target
-AD UCI Adult 48842 Sex, Race Income
-BA Bank Marketing 45221 Marital Subscription
-CA Cardio 70000 Sex Diagnosis
-CR Credit Card 30000 Sex, Education Payment
-DE Depression 11000 Sex, Race Diagnosis
-KD KDD Census 199523 Sex, Race Income
-LA Law School 20800 Sex, Race Admission
+Table 1: Experimental datasets.
+Dataset Train Validation Test Sensitive Target
+Adult 22611 11305 11306 Sex & Race Income
+Bank 22605 11303 11303 Age Subscription
+COMPAS 8322 4161 4161 Sex & Race Recidivism
 5.2 Data Processing and Baselines
-Data Processing.
-Following the data processing steps in (Kotelnikov et al., 2023), we split the numerical and categorical features for each dataset and pre-process them separately.
+5.2.1 Data Processing
+Following the data processing steps in Kotelnikov et al.
+(2023), we split the numerical and categorical features for each dataset and pre-process them separately.
 Numerical features are converted through quantile transformation, a non-parametric technique transforming variables into a standard Gaussian distribution based on quantiles, thus normalizing the data and enhancing the robustness to outliers.
 In preparation for the Markov transition process in the latent diffusion model, we encode categorical features into one-hot vectors and concatenate the resulting vectors with normalized numerical features.
-Baselines.
-In assessing the efficacy of our approach for various datasets, we conduct a comparative analysis against several baselines for tabular data synthesis.
-Within learning-based methods, we categorize them into diffusion models, generative adversarial networks, and variational autoencoders, selecting the top-performing model from each category with publicly available code.
-In addition to learning-based techniques, we incorporate SMOTE, a robust and traditional method for tabular data synthesis, into our set of baselines.
-Specifically, the selected baseline methods for comparison are described in detail as follows.
-  • TabDDPM (Kotelnikov et al., 2023): a recent diffusion-based model that has demonstrated to surpass existing methods for synthesizing tabular data.
-Leveraging multilayer perceptrons as its backbone, TabDDPM excels in learning unbalanced distributions present in the training data.
-  • CTGAN (Xu et al., 2019): a tabular data synthesis approach based on generative adversarial networks, providing publicly available code that is accessible and user-friendly in (Patki et al., 2016).
-  • TVAE (Xu et al., 2019): a state-of-the-art variational autoencoder specifically designed to generate tabular data with a user-friendly implementation in (Patki et al., 2016).
+5.2.2 Baselines
+To evaluate the effectiveness of our approach across different datasets, we perform a comparative analysis against state-of-the-art (SOTA) baseline models that have publicly available code for tabular data synthesis.
+Additionally, we compare our method with fair tabular generative models based on GAN and SMOTE.
+The specific baseline methods chosen for comparison are detailed below.
+  • Fair Class Balancing (FairCB) (Yan et al., 2020): a SMOTE-based oversampling method to adjust class distributions in the training data to address class imbalances.
+  • TabFairGAN (FairTGAN) (Rajabi & Garibay, 2022): a fairness-aware GAN tailored for tabular data incorporates a fair discriminator to ensure equality in the joint probability distribution of sensitive features and labels.
+  • CoDi (Lee et al., 2023): a recent diffusion-based approach that combines two diffusion models: one for continuous variables and another for discrete variables.
+These models are trained separately to effectively learn the distributions of their respective data types.
+However, they are also linked together by conditioning each model’s sample generation on the other model’s samples at each time step, which enhances the learning of correlations between continuous and discrete variables.
+Additionally, to further strengthen the connection between the two models, contrastive learning is applied to both diffusion models.
+Negative samples are created by breaking the inter-correlation between the continuous and discrete variable sets, thus reinforcing the models’ ability to learn these relationships.
+  • GReaT (Borisov et al., 2022): utilizes pre-trained transformer-decoder language models (LLMs) to generate synthetic tabular data through a two-stage process.
+In the first stage, it transforms each row of the real tabular dataset into a textual representation by representing each feature with its name and value in a subject-predicate-object structure.
+The transformation of these features is then concatenated in a random order to form the feature vector of the corresponding row.
+The random ordering enables the LLM to be fine-tuned on samples without depending on the sequence of features, allowing for arbitrary conditioning during tabular data generation.
+This textual representation of the tabular dataset is used to fine-tune the pre-trained LLM.
+In the second stage, the fine-tuned LLM generates synthetic data by initializing it with any combination of feature names or feature name-value pairs, allowing the model to complete the remaining feature vector in its textual representation, which can then be converted back into tabular format.
   • SMOTE (Chawla et al., 2002): a robust interpolation-based method that creates synthetic data points by combining a real data point with its $k$ nearest neighbors from the dataset.
 In our application, we leverage SMOTE to generate additional tabular data samples, employing interpolation among existing samples with the same label.
-The implementation of SMOTE is available in (Lemaître et al., 2017).
+The implementation of SMOTE is available in Lemaître et al.
+(2017).
+  • STaSy (Kim et al., 2022): a Score-based Tabular Data Synthesis (STaSy) method that adopts the score-based generative modeling approach.
+It introduces a self-paced learning method that begins by training on simpler records with a denoising score matching loss below a certain threshold.
+As training progresses and the model becomes more robust, progressively more challenging records are included in the training.
+Additionally, STaSy includes a fine-tuning approach that adjusts the model parameters by iteratively retraining it using the log probabilities of the generated samples.
+  • TabDDPM (Kotelnikov et al., 2023): a recent diffusion-based model that has demonstrated to surpass existing methods for synthesizing tabular data.
+Leveraging multilayer perceptrons as its backbone, TabDDPM excels in learning unbalanced distributions in the training data.
+  • TabSyn (Zhang et al., 2023): a SOTA method that utilizes a latent diffusion model to generate tabular data containing both numerical and categorical features.
+First, a VAE model, specifically designed for tabular-structured data, converts raw tabular data with both numerical and categorical features into embeddings in a regularized latent space that has a distribution close to the standard normal distribution.
+Following this, a score-based diffusion model is trained in the embedding space to capture the distribution of these latent embeddings.
+The diffusion model adds Gaussian noise that increases linearly over time in the forward process, which minimizes approximation errors in the reverse process, thus allowing for synthetic data to be generated in significantly fewer reverse steps.
 5.3 Assessing Synthetic Tabular Data
@@
 The other approach is a data-based method that operates independently of machine learning models and directly compares synthetic and real tabular data.
-Model-Based Evaluation. The computation procedure of model-based evaluation of synthetic tabular data in classification tasks involves the following steps.
-  1. 1.
-    Train a generative model on the original training set.
-  2. 2.
-    Generate synthetic data with the generative model.
-  3. 3.
-    Train classifiers on the synthetic data.
-  4. 4.
-    Evaluate classifiers on the original validation data.
-Specifically, when assessing model accuracy on the original validation data, it is referred to as machine learning efficiency as used in (Choi et al., 2017).
+5.3.1 Data-Based Evaluation
+We evaluate synthetic tabular data by directly comparing it with real data, with a focus on column-wise density estimation and pair-wise column correlations.
+To address privacy concerns when sharing tabular data publicly, synthetic data should not replicate the original data.
+As suggested by (Zhang et al., 2023), we compute “distance to the closest record” (DCR) scores to ensure this.
+Furthermore, when assessing the fairness of synthetic tabular data using data-based methods, we analyze the class imbalances within sensitive features like sex or race, as well as the joint distribution with the target label.
+5.3.2 Model-Based Evaluation
+The model-based evaluation process for synthetic tabular data in classification tasks involves the following steps: first, training a generative model on the original training set; next, generating synthetic data using the trained generative model; then, training classifiers on the synthetic data; and finally, evaluating these classifiers on the original test data.
+Specifically, when assessing model accuracy on the original test data, it is referred to as machine learning efficiency as used in (Choi et al., 2017).
 Beyond accuracy assessments, the evaluation of synthetic data can extend to fairness scores of classifiers, providing insights into the fairness of the generated data.
@@
   • Equal allocation, a fundamental principle in model-based fairness evaluations, entails the idea that a model should distribute resources or opportunities proportionally among different groups, irrespective of their affiliation with any privileged group.
-Equal allocation is quantified using the demographic parity ratio.
+Equal allocation is quantified using the demographic parity ratio (DPR).
 This ratio reveals the extent of the imbalance by comparing the lowest and highest selection rates (the proportion of examples predicted as positive) between groups.
@@
 This entails maintaining the same precision level for each group.
-The equalized odds ratio, which represents the smaller of two ratios comparing true and false positive rates between groups, serves as a metric for assessing performance fairness in this context.
-The evaluation of the demographic parity ratio and equalized odds ratio depends on the selection of sensitive features.
-When evaluating fairness scores in our experiments, we choose the first sensitive feature specified in Table 1.
+The equalized odds ratio (EOR), which represents the smaller of two ratios comparing true and false positive rates between groups, serves as a metric for assessing performance fairness in this context.
 Both the demographic parity ratio and equalized odds ratio are within the range $[0,1]$.
@@
 Higher values in either metric indicate progress toward fairness.
-However, attaining perfect fairness, represented by a score of $1$, may not always be feasible or appropriate depending on the specific context.
-Our proposed method is expected to have enhanced fairness scores compared to baseline models, and we interpret the results using SHAP feature importance (Lundberg & Lee, 2017).
-Data-Based Evaluation.
-The data-based evaluation directly compares synthetic and real tabular data.
-To address the privacy concerns when sharing tabular data publicly, the synthetic tabular data is expected not to be a “copy” of the original data.
-The originality of synthetic data is quantified by utilizing a nearest-neighbor distance metric suggested by (Park et al., 2018), known as the “distance to the closest record” (DCR), calculated as the Euclidean distance between each synthetic sample and its nearest real data point.
-Larger DCR values indicate that the synthetic data has a greater dissimilarity to the original data.
-In addition, when evaluating the fairness of synthetic tabular data with a data-based method, we examine the imbalance in class characteristics within sensitive features such as sex or race.
+In our experiments, when evaluating the fairness scores of classifiers as shown in Table 4, we evaluate the fairness scores of classifiers by selecting sex as the sensitive attribute for the Adult and COMPAS datasets, and age group for the Bank dataset.
 5.4 Computational Results
-This section presents the computational results of our evaluation of synthetic tabular data generated with five different random seeds.
-To ensure the robustness of our model-based evaluations, we employed a diverse set of classic classifiers, including decision trees, random forests, support vector machines, multilayer perceptrons, and XGBoost.
-We determined the optimal model for each synthetic dataset through meticulous training and evaluation across separate training and validation sets, utilizing five random seeds for each model.
-The averaged validation accuracy, along with the standard deviation, as shown in Table 2, revealed XGBoost as the dominant performer across most datasets, with random forest excelling in one exceptional case (see fairness scores in Section A.3).
-Table 2: The best and average values of validation accuracy on experimental datasets (the averaged validation accuracy $\pm$ the standard deviation)
-Dataset Model Choice Best Average
-AD XGBoost $0.866_{\pm 0.000}$ $0.795_{\pm 0.041}$
-BA XGBoost ${0.900_{\pm 0.000}}$ $0.869_{\pm 0.031}$
-CA XGBoost ${0.733_{\pm 0.000}}$ $0.658_{\pm 0.015}$
-CR Random Forest ${0.816_{\pm 0.000}}$ $0.757_{\pm 0.018}$
-DE XGBoost ${0.941_{\pm 0.000}}$ $0.929_{\pm 0.001}$
-KD XGBoost ${0.956_{\pm 0.000}}$ $0.946_{\pm 0.000}$
-LA XGBoost ${1.000_{\pm 0.000}}$ $0.942_{\pm 0.004}$
-Performance.
-In our exploration of the machine learning efficiency of synthetic data compared to baselines, we employ two rigorous protocols.
-The first protocol calculates the average machine learning efficiency across all classifiers, while the second identifies and utilizes the best classifier for each dataset.
-This multifaceted approach ensures a robust and insightful evaluation.
-As illustrated in Table 3, our proposed method maintains competitiveness in terms of average machine learning efficiency, even when incorporating a uniformly distributed direction for sensitive features during the sampling phase.
-It generally outperforms CTGAN and TVAE.
-Moreover, our method excels against TabDDPM on the majority of datasets.
-While the interpolation-based method, SMOTE, proves highly competitive and challenging to outperform, our model achieves a machine learning efficiency very close to SMOTE, underscoring its effectiveness.
-The best machine learning efficiency exhibits a similar pattern with the average machine learning efficiency, with SMOTE and our model emerging as the two most competitive methods.
-Table 3: The values of machine learning efficiency averaged on all types of classifiers
-Average Accuracy
-AD BA CA CR DE KD LA
-TabDDPM $0.777_{\pm 0.093}$ $0.874_{\pm 0.017}$ $0.640_{\pm 0.025}$ $0.768_{\pm 0.032}$ $0.920_{\pm 0.001}$ $0.937_{\pm 0.001}$ $0.905_{\pm 0.012}$
-CTGAN $0.742_{\pm 0.148}$ $0.873_{\pm 0.019}$ $0.600_{\pm 0.061}$ $0.746_{\pm 0.058}$ $0.867_{\pm 0.050}$ $0.931_{\pm 0.014}$ $0.882_{\pm 0.112}$
-TVAE $0.786_{\pm 0.023}$ $0.854_{\pm 0.026}$ $0.613_{\pm 0.058}$ $0.784_{\pm 0.021}$ $0.881_{\pm 0.051}$ $0.938_{\pm 0.008}$ $0.919_{\pm 0.119}$
-SMOTE $0.807_{\pm 0.038}$ $0.880_{\pm 0.022}$ $0.632_{\pm 0.074}$ $0.770_{\pm 0.044}$ $0.928_{\pm 0.022}$ $0.939_{\pm 0.013}$ $0.930_{\pm 0.124}$
-Ours $0.803_{\pm 0.022}$ $0.880_{\pm 0.011}$ $0.629_{\pm 0.001}$ $0.734_{\pm 0.045}$ $0.926_{\pm 0.004}$ $0.932_{\pm 0.004}$ $0.921_{\pm 0.003}$
-Best Accuracy
-AD BA CA CR DE KD LA
-TabDDPM $0.834_{\pm 0.004}$ $0.895_{\pm 0.003}$ $0.708_{\pm 0.003}$ $0.816_{\pm 0.002}$ $0.937_{\pm 0.003}$ $0.945_{\pm 0.001}$ $0.964_{\pm 0.002}$
-CTGAN $0.817_{\pm 0.002}$ ${0.888}_{\pm 0.002}$ ${0.652}_{\pm 0.006}$ ${0.716}_{\pm 0.009}$ ${0.870}_{\pm 0.008}$ ${0.941}_{\pm 0.001}$ ${0.975}_{\pm 0.002}$
-TVAE ${0.810}_{\pm 0.004}$ ${0.849}_{\pm 0.008}$ ${0.662}_{\pm 0.006}$ ${0.773}_{\pm 0.006}$ ${0.900}_{\pm 0.003}$ ${0.944}_{\pm 0.001}$ ${0.993}_{\pm 0.001}$
-SMOTE $0.853_{\pm 0.000}$ $0.898_{\pm 0.000}$ $0.704_{\pm 0.000}$ $0.817_{\pm 0.001}$ $0.941_{\pm 0.000}$ $0.949_{\pm 0.000}$ $1.000_{\pm 0.000}$
-Ours $0.838_{\pm 0.001}$ $0.892_{\pm 0.003}$ $0.709_{\pm 0.004}$ $0.806_{\pm 0.004}$ $0.942_{\pm 0.003}$ $0.930_{\pm 0.003}$ $0.995_{\pm 0.004}$
-Privacy.
-As explained in Section 5.3, we employ DCR values to assess the privacy score of the synthetic tabular data.
-The median DCR values, obtained across five random seeds, are detailed in Table 4.
-Remarkably, our method consistently attains a superior privacy score in comparison to both SMOTE and TabDDPM.
-This suggests that our approach excels in maintaining the distinctiveness and privacy of the synthetic data.
-Table 4: The median value of distance to the closest record
-AD BA CA CR DE KD LA
-TabDDPM $0.179$ $0.195$ $0.021$ $0.123$ $0.520$ $2.089$ $0.176$
-CTGAN $0.340$ $0.303$ $0.028$ $0.202$ $0.428$ $2.451$ $0.276$
-TVAE $0.253$ $1.415$ $0.032$ $0.208$ $0.443$ $2.018$ $1.414$
-SMOTE $0.170$ $0.191$ $0.021$ $0.117$ $0.398$ $2.005$ $0.178$
-Ours $0.511$ $0.229$ $0.023$ $0.255$ $0.430$ $2.831$ $0.216$
-Fairness.
-A key contribution of our work is the generation of balanced tabular data while considering sensitive features.
-To substantiate our claim and provide empirical evidence, we conduct a thorough assessment of the fairness of our proposed model.
-After choosing the best-performing classifier for each dataset, we assess its fairness by evaluating two metrics: demographic parity ratio and equalized odds ratio, and the results are presented in Table 5.
-Our approach outperforms baseline methods notably in terms of equal allocation, as evidenced by higher values of demographic parity ratio across nearly all datasets.
-For instance, on the KDD Census dataset, our method achieves a demographic parity ratio of $0.613$, surpassing baseline methods like SMOTE and TabDDPM by around fifty percent, while SMOTE only attains $0.141$, and TabDDPM reaches $0.097$.
-Similarly, on the UCI Adult dataset, our model demonstrates a demographic parity ratio of $0.529$, surpassing the ratios of $0.306$ by SMOTE and $0.312$ by TabDDPM.
-In addition, our model demonstrates competitiveness in terms of equal performance metrics, as indicated by the equalized odds ratio.
-Particularly noteworthy are datasets such as UCI Adult and KDD Census, where our model achieves an equalized odds ratio of $0.641$ and $0.884$, respectively, surpassing baseline models that register less than $0.250$ in these metrics.
-Furthermore, employing equitable synthetic data to train classifiers can yield higher fairness scores compared to training with the original data, as illustrated in Table 7.
-In contrast, baseline models generally achieve comparable fairness scores.
-Table 5: The values of demographic parity ratio and equalized odds ratio with the best-performing classifier
-Demographic Parity Ratio
-AD BA CA CR DE KD LA
-TabDDPM $0.312_{\pm 0.033}$ $0.590_{\pm 0.070}$ $0.957_{\pm 0.036}$ $0.791_{\pm 0.063}$ $0.634_{\pm 0.020}$ $0.097_{\pm 0.032}$ $0.858_{\pm 0.007}$
-TabGAN ${0.188}_{\pm 0.039}$ ${0.749}_{\pm 0.060}$ ${0.760}_{\pm 0.027}$ ${0.915}_{\pm 0.051}$ ${0.706}_{\pm 0.021}$ ${0.112}_{\pm 0.037}$ ${0.844}_{\pm 0.006}$
-TabVAE ${0.249}_{\pm 0.017}$ ${0.624}_{\pm 0.045}$ ${0.965}_{\pm 0.011}$ ${0.740}_{\pm 0.088}$ ${0.664}_{\pm 0.012}$ ${0.092}_{\pm 0.027}$ ${0.847}_{\pm 0.006}$
-SMOTE ${0.306}_{\pm 0.000}$ ${0.625}_{\pm 0.000}$ ${0.975}_{\pm 0.000}$ ${0.787}_{\pm 0.000}$ ${0.671}_{\pm 0.000}$ ${0.141}_{\pm 0.000}$ ${0.851}_{\pm 0.000}$
-Ours $0.529_{\pm 0.046}$ $0.839_{\pm 0.098}$ $0.963_{\pm 0.017}$ $0.828_{\pm 0.048}$ $0.706_{\pm 0.005}$ $0.613_{\pm 0.057}$ $0.862_{\pm 0.007}$
-Equalized Odds Ratio
-AD BA CA CR DE KD LA
-TabDDPM $0.247_{\pm.0.033}$ $0.496_{\pm.0.077}$ $0.919_{\pm.0.068}$ $0.680_{\pm.0.081}$ $0.764_{\pm.0.110}$ $0.134_{\pm.0.050}$ $0.593_{\pm.0.484}$
-CTGAN ${0.125}_{\pm 0.047}$ ${0.760}_{\pm 0.074}$ ${0.669}_{\pm 0.038}$ ${0.895}_{\pm 0.048}$ ${0.815}_{\pm 0.062}$ ${0.178}_{\pm 0.067}$ ${0.611}_{\pm 0.172}$
-TVAE ${0.205}_{\pm 0.023}$ ${0.621}_{\pm 0.046}$ $0.950_{\pm 0.027}$ ${0.667}_{\pm 0.115}$ ${0.825}_{\pm 0.067}$ ${0.126}_{\pm 0.046}$ ${0.562}_{\pm 0.175}$
-SMOTE ${0.231}_{\pm 0.000}$ ${0.593}_{\pm 0.000}$ ${0.939}_{\pm 0.000}$ ${0.683}_{\pm 0.000}$ ${0.827}_{\pm 0.000}$ ${0.202}_{\pm 0.000}$ ${1.000}_{\pm 0.000}$
-Ours $0.641_{\pm 0.108}$ $0.760_{\pm 0.043}$ $0.901_{\pm 0.035}$ $0.802_{\pm 0.056}$ $0.254_{\pm 0.089}$ $0.884_{\pm 0.055}$ $0.170_{\pm 0.340}$
-Figure 2: Comparison in SHAP values on KDD dataset: balanced scenario (top) vs. unbalanced scenario (bottom) using XGBoost
-In Figure 2, we illustrate the feature importance through absolute SHAP values when making predictions with XGBoost on the KDD Census dataset.
-On the top, we present the scenario where synthetic data generated by our approach exhibits a uniformly distributed sensitive feature, sex, across all samples.
-In contrast, the figure on the bottom showcases data generated by TabDDPM with biased sampling in the sex feature.
-It is evident that in the data generated by TabDDPM, the variable sex significantly influences decision-making, while our method mitigates the impact of sex on the final decision to a considerable extent.
-Apart from fairness evaluation dependent on machine learning models, we also analyze the class distribution within the sensitive attributes.
-For example, in the UCI Adult dataset, the distribution of sensitive attributes has an inherent bias.
-In contrast, the synthetic data generated through our approach exhibits more balanced distributions, as depicted in Figure 3, where the y-axis represents the percentage of the respective attribute.
-We present visualizations of class imbalance in sensitive attributes for the rest of experimental datasets in Figure 4.
-Figure 3: Comparison in the real vs. synthetic distribution of sensitive attributes in the UCI Adult dataset
-5.5 Limitations and Discussion
-Our proposed method is designed to generate balanced synthetic tabular data considering sensitive attributes.
-Empirically, classifiers trained on such balanced synthetic tabular data demonstrate higher fairness scores at equal allocation metrics indicated by demographic parity ratios.
-However, there is no guarantee that equal performance, reflected by equalized odds ratios, will always improve.
-We use U-Net with attention layers as the backbone neural network for the posterior estimator in the diffusion framework, but it is time-consuming.
-Despite our model exhibiting superior machine learning efficiency and higher fairness scores compared to other deep generative models in our experiments, the computational time is approximately five times longer.
-In the future, we may explore the incorporation of constraints to ensure equal performance and seek ways to reduce computational costs.
+This section presents the computational results of our evaluation of synthetic tabular data generated with three different random seeds.
+We trained the SOTA machine learning model for tabular data, CatBoost, on three different sets of synthetic data, each generated with a unique random seed.
+5.4.1 Fidelity and Diversity of Synthetic Data
+Table 2: The values of error rates of column-wise density estimation and pair-wise column correlations.
+Adult Bank COMPAS Mean Adult Bank COMPAS Mean
+[…]
+CoDi .$145_{\pm.000}.$$154_{\pm.000}.$$205_{\pm.001}.$ $16.8\%$$495_{\pm.000}.$$344_{\pm.001}.$$550_{\pm.001}.$ $46.3\%$
+GReaT .$077_{\pm.000}.$$102_{\pm.001}.$$093_{\pm.001}.$ $9.1\%$$208_{\pm.015}.$$213_{\pm.010}.$$165_{\pm.016}.$ $19.5\%$
+SMOTE .$025_{\pm.000}.$$020_{\pm.000}.$$021_{\pm.001}.$ $2.2\%$$054_{\pm.004}.$$042_{\pm.005}.$$047_{\pm.005}.$ $4.8\%$
+STaSy .$102_{\pm.000}.$$182_{\pm.000}.$$108_{\pm.001}.$ $13.1\%$$163_{\pm.000}.$$221_{\pm.001}.$$138_{\pm.001}.$ $17.4\%$
+TabDDPM .$037_{\pm.001}.$$028_{\pm.001}.$$057_{\pm.001}.$ $4.1\%$$055_{\pm.001}.$$052_{\pm.001}.$$090_{\pm.001}.$ $6.6\%$
+TabSyn .$010_{\pm.000}.$$009_{\pm.000}.$$027_{\pm.001}.$ $1.5\%$$035_{\pm.001}.$$033_{\pm.002}.$$054_{\pm.001}.$ $4.1\%$
+FairCB .$076_{\pm.000}.$$066_{\pm.000}.$$039_{\pm.000}.$ $6.0\%$$125_{\pm.000}.$$111_{\pm.000}.$$074_{\pm.000}.$ $10.3\%$
+FairTGAN .$034_{\pm.000}.$$030_{\pm.000}.$$055_{\pm.001}.$ $4.0\%$$080_{\pm.001}.$$053_{\pm.000}.$$087_{\pm.001}.$ $7.3\%$
+Ours .$126_{\pm.001}.$$121_{\pm.000}.$$109_{\pm.001}.$ $11.9\%$$201_{\pm.000}.$$174_{\pm.005}.$$174_{\pm.003}.$ $18.3\%$
+Table 3: The values of machine learning efficiency with the best classifier for each dataset and DCR scores.
+Adult Bank COMPAS Mean Adult Bank COMPAS Mean
+[…]
+Real .$928_{\pm.000}.$$936_{\pm.000}.$$810_{\pm.001}.$ $89.1\%$
+CoDi .$858_{\pm.001}.$$826_{\pm.014}.$$678_{\pm.002}.$ $78.7\%$$331_{\pm.003}.$$348_{\pm.001}.$$400_{\pm.002}.$ $36.0\%$
+GReaT .$901_{\pm.002}.$$688_{\pm.024}.$$717_{\pm.008}.$ $76.9\%$$320_{\pm.002}.$$348_{\pm.003}.$$377_{\pm.003}.$ $34.8\%$
+SMOTE .$914_{\pm.000}.$$928_{\pm.001}.$$778_{\pm.002}.$ $87.3\%$$327_{\pm.004}.$$265_{\pm.001}.$$273_{\pm.003}.$ $28.8\%$
+STaSy .$885_{\pm.003}.$$895_{\pm.003}.$$728_{\pm.013}.$ $83.6\%$$344_{\pm.001}.$$345_{\pm.002}.$$362_{\pm.001}.$ $35.0\%$
+TabDDPM .$907_{\pm.001}.$$917_{\pm.002}.$$745_{\pm.001}.$ $85.6\%$$339_{\pm.000}.$$350_{\pm.002}.$$367_{\pm.005}.$ $35.2\%$
+TabSyn .$911_{\pm.000}.$$919_{\pm.000}.$$749_{\pm.001}.$ $86.0\%$$339_{\pm.002}.$$351_{\pm.005}.$$367_{\pm.006}.$ $35.2\%$
+FairCB .$915_{\pm.001}.$$907_{\pm.002}.$$771_{\pm.001}.$ $86.4\%$$054_{\pm.000}.$$031_{\pm.001}.$$012_{\pm.001}.$ $3.2\%$
+FairTGAN .$881_{\pm.000}.$$863_{\pm.006}.$$705_{\pm.002}.$ $81.6\%$$348_{\pm.002}.$$348_{\pm.004}.$$374_{\pm.006}.$ $35.7\%$
+Ours .$893_{\pm.002}.$$914_{\pm.002}.$$734_{\pm.001}.$ $84.7\%$$344_{\pm.001}.$$350_{\pm.002}.$$370_{\pm.004}.$ $35.5\%$
+Table 2 and Table 3 show that our method remains competitive with SOTA models in terms of column-wise density estimation, pair-wise column correlations, and machine learning efficiency, even after incorporating a uniformly distributed direction for sensitive features during the sampling phase.
+Across the three experimental datasets, our model achieved an average AUC of $84.7\%$, which is only $1.3\%$ lower than TabSyn, the SOTA deep learning tabular synthesis method.
+Although altering the original sensitive distribution is anticipated to reduce AUC, the results demonstrate that our method effectively preserves high fidelity in the generated synthetic data.
+While SMOTE and FairCB excel in fidelity, their low DCR scores suggest that they closely mimic the real data, which is expected given their direct interpolation of the original data points.
+5.4.2 Fairness Scores in Classification Tasks
+Table 4: The values of DPR and EOR with best classifier for each dataset.
+Adult Bank COMPAS Mean Adult Bank COMPAS Mean
+[…]
+Real .$309_{\pm.001}.$$402_{\pm.015}.$$675_{\pm.006}.$ $46.2\%$$193_{\pm.005}.$$367_{\pm.024}.$$645_{\pm.015}.$ $40.2\%$
+CoDi .$293_{\pm.034}.$$189_{\pm.054}.$$855_{\pm.025}.$ $44.6\%$$247_{\pm.042}.$$172_{\pm.063}.$$857_{\pm.031}.$ $42.5\%$
+GReaT .$249_{\pm.015}.$$572_{\pm.289}.$$624_{\pm.066}.$ $48.2\%$$155_{\pm.028}.$$380_{\pm.126}.$$543_{\pm.075}.$ $35.9\%$
+SMOTE .$321_{\pm.006}.$$405_{\pm.028}.$$648_{\pm.021}.$ $45.8\%$$254_{\pm.011}.$$381_{\pm.026}.$$589_{\pm.009}.$ $40.8\%$
+STaSy .$261_{\pm.045}.$$468_{\pm.123}.$$436_{\pm.092}.$ $38.8\%$$182_{\pm.069}.$$451_{\pm.113}.$$433_{\pm.140}.$ $35.5\%$
+TabDDPM .$261_{\pm.006}.$$337_{\pm.020}.$$558_{\pm.041}.$ $38.5\%$$156_{\pm.007}.$$334_{\pm.028}.$$540_{\pm.047}.$ $34.3\%$
+TabSyn .$281_{\pm.017}.$$336_{\pm.016}.$$697_{\pm.040}.$ $43.8\%$$178_{\pm.019}.$$317_{\pm.023}.$$664_{\pm.062}.$ $38.6\%$
+FairCB .$286_{\pm.002}.$$719_{\pm.005}.$$675_{\pm.006}.$ $56.0\%$$192_{\pm.003}.$$801_{\pm.016}.$$638_{\pm.021}.$ $54.4\%$
+FairTGAN .$554_{\pm.006}.$$338_{\pm.093}.$$448_{\pm.025}.$ $44.7\%$$697_{\pm.017}.$$158_{\pm.033}.$$392_{\pm.041}.$ $41.6\%$
+Ours .$543_{\pm.026}.$$710_{\pm.057}.$$800_{\pm.080}.$ $68.4\%$$667_{\pm.059}.$$649_{\pm.040}.$$825_{\pm.113}.$ $71.4\%$
+As shown in Table 4, our method significantly outperforms all baseline models, achieving a mean DPR of $68.4\%$ and a mean EOR of $71.4\%$.
+In comparison, state-of-the-art tabular generative models and FairTGAN fall below $50\%$ for both fairness metrics, while FairCB remains below $60\%$ for both.
+From Table 3, our method achieves an average AUC of $84.7$, which is higher than FairTGAN and only $1.6$ lower than FairCB.
+By jointly optimizing performance and fairness, our method presents a more balanced trade-off than existing approaches.
+Moreover, in Section E.1, we demonstrate that simply overwriting the distribution of sensitive attributes is insufficient to mitigate bias in downstream classification tasks and can lead to a decrease in classification performance.
+5.4.3 Sensitive Feature Distributions
+In addition to evaluating the fairness of machine learning models, we also examine the class distribution of sensitive attributes in synthetic data.
+To achieve this, we use stacked bar plots to visualize the contingency tables for sensitive features and the target label.
+We further compare the distribution of sensitive attributes between real and synthetic data using bar plots.
+Figure 2: The distribution of sensitive attributes across different target label values on the Adult dataset using our method.
+We present the distribution of sensitive attributes versus the target label on the Adult dataset using our method in Figure 2.
+The synthetic data generated by our approach demonstrates a balanced joint distribution between sensitive features and the target label, effectively reducing the disparities observed in the original data.
+Additionally, we provide plots of the real data distribution and synthetic distributions generated by other methods in Appendix D.
+Note that we only visualize the synthetic data generated by our method, FairTGAN, and FairCB, which are three fairness-aware generative models, because we assume that the other baseline models replicate the real data distribution.
+In comparison, our method effectively produces a more balanced joint distribution of sensitive attributes and target labels than the FairTGAN, FairCB, and the real distribution.
+Figure 3: Comparison in the real versus synthetic distribution of sensitive attributes across all datasets.
+5.4.4 Performance-Fairness Trade-Off
+Balancing fairness and predictive accuracy is a key challenge in synthetic data generation, particularly when fairness improvements might impact important outcomes.
+This section explores how adjustments in fairness levels influence both performance and fairness metrics in the generated data.
+Our method provides a way to control this trade-off, allows practitioners to find a balance that meets their specific needs, whether they prioritize fairness, accuracy, or a compromise between the two.
+To analyze the trade-off between fairness and accuracy, we define a balancing level $i$ ranging from 00 to $10$.
+This level quantifies how much we adjust the original joint distribution of the target and sensitive attributes to make it more uniform.
+Suppose we have $N$ sensitive attributes, each attribute with $C_{i}$ unique values, where $i$ ranges from 1 to $N$.
+This results in $K$ combinations of values where $K=C_{i}\times\cdots\times C_{N}$.
+Under the same targe label, each combination has a probability $y_{k}$ where $k$ is one of the $K$ combinations.
+For each combination, we calculate an offset $d_{k}=\bar{y}-y_{k}$, the difference between the average probability $\bar{y}={\sum_{k}y_{k}}/{|\mathbf{y}|}$ and the current probability $y_{k}$.
+The offset $d_{k}$ represents how much the current joint probability deviation $y_{k}$ for a combination of sensitive attributes and the target label deviate from the average probability $\bar{y}$.
+By identifying these deviations, we can measure how far the joint distribution is from a uniform distribution, which is often desired for achieving fairness.
+For a balancing level $i\in[0,10]$, the balancing ratio is $i/10$, and the adjusted values $y_{k}^{\text{balanced}}$ are calculated as:
+$y_{k}^{\text{balanced}}=y_{k}+d_{k}\times\frac{i}{10}$ (8)
+The balanced values $y_{k}^{\text{balanced}}$ represent the adjusted probabilities for each combination of sensitive attributes and the target label after applying a fairness transformation.
+The balancing process ensures that $y_{k}^{\text{balanced}}$ systematically reduces disparities in the joint distribution of sensitive attributes and target labels, moving it closer to uniformity as the balancing level $i$ increases.
+By choosing an appropriate balancing level, the method allows for a practical balance between performance and fairness.
+The results presented in the previous sections are based on a balancing level of 10, which achieves a nearly uniform joint distribution of the target and sensitive attributes.
+We conducted experiments to assess how different balancing levels affect the performance-fairness trade-off.
+We evaluated metrics such as AUC, DPR, and EOR, with DPR and EOR averaged across all sensitive attributes.
+We computed a composite score as a weighted sum of these metrics, with weights of $0.5$ for AUC, $0.25$ for DPR, and $0.25$ for EOR.
+Figure 4 illustrates the trade-off between fairness and performance across different balancing levels.
+The optimal composite scores for the Adult and Bank datasets are achieved at a balancing level of $10$, indicating a uniform distribution.
+For the COMPAS dataset, the best score is at level $9$, representing a nearly uniform distribution.
+Practitioners can adjust the metric weights to suit specific application needs or establish thresholds for individual metrics based on their requirements.
+Figure 4: Trade-off between performance and fairness across balancing levels for experimental datasets.
+SCORE was computed as a weighted sum of these metrics, with weights of 0.5 for AUC, 0.25 for DPR, and 0.25 for EOR.
+The best composite scores for the Adult and Bank datasets are achieved at a balancing level of 10, while the COMPAS dataset achieves the best score at level 9.
+6 Limitations and Discussion
+Our proposed method is designed to generate balanced synthetic tabular data while considering sensitive attributes.
+However, it requires sensitive features to be specified in advance, which can be challenging in large-scale enterprise datasets with thousands of features.
+Furthermore, we could extend our testing to datasets with more than two sensitive attributes and imbalanced target distributions.
+In terms of efficiency, we use a U-Net with attention layers as the backbone neural network for the posterior estimator in the diffusion framework, but this approach is more time-consuming compared to TabSyn.
+As the number of sensitive attributes increases, there are more hyperparameters to tune, which makes fine-tuning more computationally demanding, especially with the long sampling time.
+In the future, we aim to explore methods to reduce computational costs.
+We believe the main bottlenecks in computational costs are due to the long diffusion timesteps and the quadratic complexity of attention modules.
+Possible solutions include reducing diffusion timesteps via acceleration methods such as Flow-DPM-Solver Xie et al.
+(2024) and refining the backbone, for instance, by employing linear attention Katharopoulos et al.
+(2020) or by using a pre-trained autoencoder to embed tabular data into a lower-dimensional latent space.
+Lastly, our proposed method is designed to condition on specified group labels, which limits its applicability in scenarios where demographic information is unavailable or restricted due to privacy concerns.
+Future iterations of the model could integrate privacy-preserving techniques, such as synthetic attribute imputation, to enable fairness-aware data generation without requiring explicit sensitive attribute disclosure.
+On the other hand, our proposed method generates fair data with a balanced joint distribution of the target label and multiple sensitive attributes, achieving higher fairness scores than baseline methods.
+Additionally, the model shows strong potential for extending to multimodal data synthesis.
+If we can improve the efficiency of our model and successfully adapt it to multimodal scenarios, we believe it could be widely applied to address fairness issues in fields such as healthcare, finance, and beyond.
 7 Conclusion
@@
 An adverse possibility is that bad individuals could manipulate distributions in sensitive attributes to generate biased (even stronger than original) data to harm minority groups.
+Additionally, the widespread release of synthetic data can diminish the quality of data sources available on the internet.
+Repeatedly training generative models on synthetic data in a self-consuming loop can lead to a decline in the quality of data synthesis, as discussed in (Alemohammad et al., 2023).
 Appendix A Details about Datasets
-A.1 Data Sources
-A.1.1 UCI Adult
-The UCI Adult dataset (Kohavi et al., 1996) is widely used as a benchmark for exploring fairness and bias in machine learning.
+A.1 Adult
+The Adult dataset (Kohavi et al., 1996) is widely used as a benchmark for exploring fairness and bias in machine learning.
 It contains 48842 data points.
@@
 Download it from OpenML.
-A.1.2 Bank Marketing
-The Bank Marketing dataset is collected from direct marketing campaigns of a Portuguese banking institution.
+A.2 Bank Marketing
+The Bank Marketing dataset (Moro et al., 2014) is collected from direct marketing campaigns of a Portuguese banking institution.
 Comprising 45211 instances with 16 features, the predicted outcome is a binary variable indicating whether a client subscribed to a term deposit.
@@
 Download it from OpenML.
-A.1.3 Cardio
-The Cardio dataset focuses on diagnosing cardiovascular disease (yes or no) and it contains 70000 instances.
-It incorporates 11 features, which contain demographics and factual information (blood pressure, cholesterol levels, etc.), as well as self-reported labels such as whether smoking or not.
-The sensitive feature is sex.
+A.3 COMPAS
+The COMPAS (Correctional Offender Management Profiling for Alternative Sanctions) dataset (Larson et al., 2016) is derived from a risk assessment tool used in the criminal justice system to evaluate the likelihood of recidivism among criminal defendants.
+It includes data on individuals arrested in Broward County, Florida, with 8 features and a binary outcome indicating whether the individual was predicted to reoffend.
+Features encompass demographic information, criminal history, and COMPAS risk scores, with sex and race being sensitive attributes.
 Download it from OpenML.
-A.1.4 Credit Card
-The Credit Card dataset (Yeh & Lien, 2009) explores customer default payments (yes or no) among credit card clients in Taiwan, and it contains 30000 data points.
-Each data point has 23 features including the information of demographics and financial records (monthly payments, etc.).
-The sensitive features are sex and education.
-Download it from OpenML.
-A.1.5 Depression
-The Depression dataset contains demographic and clinical data (concentration, rumination, etc.) to predict a diagnosis of depression.
-It has 11000 instances, each with 18 features, and sensitive features are sex and race.
-Download it from OpenML.
-A.1.6 KDD Census
-The KDD Census dataset is an expanded version of the UCI Adult dataset.
-It involves 199523 instances, each featuring 41 attributes and a binary target variable indicating income over fifty thousand dollars per year.
-The dataset covers employment, education, and demographic details, with sex and race as sensitive features.
-Download it from OpenML.
-A.1.7 Law School
-The Law School dataset presents a survey on the law school admission process.
-It has 20800 instances.
-Each instance has 12 attributes and a binary target variable indicating admission to law school.
-Demographic and academic information, such as GPA, are included in features.
-The sensitive features are sex and race.
-Download it from OpenML.
-A.2 Number of Features
-Dataset Number of Numerical Features Number of Categorical Features
-UCI Adult 6 8
-Bank Marketing 7 9
-Cardio 5 6
-Credit Card 14 9
-Depression 14 4
-KDD Census 7 34
-Law School 6 6
-Table 6: The number of numerical attributes and categorical attributes in experimental datasets
-A.3 Evaluation of Best-Performing Classifiers on Original Datasets
-Dataset Model Choice Accuracy Demographic Parity Ratio Equalized Odds Ratio
-UCI Adult XGBoost $0.866$ $0.353$ $0.273$
-Bank Marketing XGBoost $0.900$ $0.619$ $0.540$
-Cardio XGBoost $0.733$ $0.992$ $0.989$
-Credit Card Random Forest $0.816$ $0.778$ $0.834$
-Depression XGBoost $0.941$ $0.696$ $0.644$
-KDD Census XGBoost $0.956$ $0.151$ $0.170$
-Law School XGBoost $1.000$ $0.825$ $1.000$
-Table 7: The values of accuracy and fairness scores of the best-performing classifier on experimental datasets
-A.4 Class Imbalance
-(a) Bank Marketing
-(b) Cardio
-(c) Credit Card
-(d) Depression
-(e) KDD Census
-(f) Law School
-Figure 4: Visualizations of class imbalance in sensitive attributes in experimental datasets except UCI Adult
+Appendix B Reproducibility
+We performed our experiments on Ubuntu 20.04.2 LTS, utilizing Python version 3.10.14.
+Our framework of choice was PyTorch 2.3.0, with CUDA version 12.1, and we ran our computations on an NVIDIA GeForce RTX 3090.
+The hyperparameters we used for our method are listed as follows.
+The variable names can be found in our attached code.
+We tune the hyperparameters using Optuna Akiba et al.
+(2019) to optimize the validation AUC.
+Table 5: The values of hyperparameters for our method.
+Notation Variable Meaning Value/Search Space
+- batch_size Batch size $256$
+$\delta$ warmup_steps Warm-up timestep 00
+$\omega_{s}$ cond_guid_weight Sensitive guidance weight $1.0$
+$\lambda$ cond_guid_threshold Threshold of the “security gate“ $1.0$
+$\omega_{m}$ cond_momentum_weight Momentum weight $0.5$
+$\beta$ cond_momentum_beta Momentum correction factor $0.5$
+$\omega_{g}$ overall_guid_weight Guidance weight $1.0$
+$T$ n_timesteps Number of timesteps in the diffusion process $\{100,1000\}$
+- n_epochs Number of training epochs $\{100,500,1000\}$
+- lr Learning rate $[0.00001,0.003]$
+For deep learning-based methods, we use the original hyperparameter settings provided in their respective implementations and mainly fine-tune the number of epochs, learning rate, and number of diffusion timesteps (if applicable) to optimize performance.
+For SMOTE-based methods, we adjust the number of nearest neighbors to achieve better results.
+We have detailed the hyperparameter search spaces for all baseline methods in Table 6.
+The hyperparameter search spaces are the same for all datasets.
+Table 6: The values of hyperparameters for baseline methods.
+Method Variable Meaning Value/Search Space
+CoDi total_epochs_both Number of training epochs $\{100,500,1000,3000\}$
+n_timesteps Number of timesteps in the diffusion process $\{100,1000\}$
+GReaT batch_size Batch size $\{4,8\}$
+n_epochs Number of training epochs $\{5,10,20\}$
+SMOTE knn Number of nearest neighbors for interpolation $\{2,21\}$
+STaSy lr Learning rate $[0.00001,0.003]$
+n_epochs Number of training epochs $\{100,500,1000\}$
+TabDDPM lr Learning rate $[0.00001,0.003]$
+n_epochs Number of training epochs $\{100,500,1000\}$
+num_timesteps Number of timesteps in the diffusion process $\{100,1000\}$
+TabSyn lr Learning rate $[0.001,0.002]$
+n_epochs Number of training epochs $\{4000\}$
+FairCB knn Number of nearest neighbors for interpolation $\{2,21\}$
+FairTGAN lr Learning rate $[0.00001,0.003]$
+n_epochs Number of training epochs $\{100,500,1000\}$
+fair_epochs Number of fairness-related training epochs $\{100,500\}$
+Appendix C Efficiency
+Table 7 provides a summary of the number of parameters, training time, and sampling time for various tabular data synthesis methods.
+Our proposed method is moderately sized, with 482,545 parameters, which is larger than TabDDPM but significantly smaller than most deep learning-based models, except FairTGAN.
+In terms of training efficiency, our method achieves a training time of 11.8 seconds per epoch, which is comparable to CoDi (10.5 seconds per epoch).
+However, the sampling time is a significant limitation, as it takes approximately 1,200 seconds to generate a synthetic dataset of the same size as the training data, making it substantially slower than all other baseline methods.
+Table 7: Comparison of training and sampling efficiency across methods
+Method Number of Parameters Training Time (s/epoch) Sampling Time (s)
+CoDi 2206711 10.5 9.1
+GReaT 81912576 725.5 236.3
+SMOTE - 51.1 17.0
+STaSy 10653326 1.4 25.1
+TabDDPM 109672 6.0 118.9
+TabSyn 10575912 0.6 2.7
+FairCB - 64.3 5.4
+FairTGAN 45368 1.9 0.3
+Ours 482545 11.8 1213.9
+Appendix D Stacked Bar Plots of Contingency Tables
+D.1 Adult
+Figure 5: Comparison of models on the Adult dataset.
+D.2 Bank Marketing
+Figure 6: Comparison of models on the Bank Marketing dataset.
+D.3 COMPAS
+Figure 7: Comparison of models on the COMPAS dataset.
+Appendix E Additional Numerical Results
+E.1 Performance and Fairness Metrics with Sensitive Features Replaced by Uniform Distribution
+We present performance and fairness of CatBoost trained on datasets with sensitive attribute replaced by uniform distribution in Table 8 and Table 9 respectively.
+Table 8: The values of machine learning efficiency with the best classifier for each dataset. The sensitive attributes are replaced by uniform distribution.
+Adult Bank COMPAS Mean
+[…]
+Real .$928_{\pm.000}.$$936_{\pm.000}.$$799_{\pm.000}.$ $88.8\%$
+CoDi .$858_{\pm.002}.$$827_{\pm.015}.$$678_{\pm.008}.$ $78.8\%$
+GReaT .$900_{\pm.003}.$$690_{\pm.026}.$$714_{\pm.010}.$ $76.8\%$
+SMOTE .$914_{\pm.000}.$$928_{\pm.001}.$$772_{\pm.002}.$ $87.1\%$
+STaSy .$883_{\pm.000}.$$894_{\pm.003}.$$722_{\pm.014}.$ $83.3\%$
+TabDDPM .$906_{\pm.001}.$$917_{\pm.002}.$$745_{\pm.002}.$ $85.6\%$
+TabSyn .$910_{\pm.000}.$$918_{\pm.001}.$$745_{\pm.001}.$ $85.8\%$
+FairCB .$914_{\pm.000}.$$907_{\pm.001}.$$765_{\pm.001}.$ $86.2\%$
+FairTGAN .$885_{\pm.002}.$$873_{\pm.006}.$$703_{\pm.002}.$ $82.0\%$
+Ours .$895_{\pm.001}.$$913_{\pm.002}.$$732_{\pm.001}.$ $84.7\%$
+Table 9: The values of DPR and EOR with best classifier for each dataset. The sensitive attributes are replaced by uniform distribution.
+Adult Bank COMPAS Mean Adult Bank COMPAS Mean
+[…]
+Real .$307_{\pm.004}.$$415_{\pm.008}.$$790_{\pm.005}.$ $50.4\%$$192_{\pm.008}.$$368_{\pm.008}.$$857_{\pm.007}.$ $47.2\%$
+CoDi .$345_{\pm.030}.$$402_{\pm.063}.$$901_{\pm.011}.$ $54.9\%$$328_{\pm.038}.$$556_{\pm.125}.$$923_{\pm.010}.$ $60.2\%$
+GReaT .$256_{\pm.010}.$$622_{\pm.269}.$$814_{\pm.024}.$ $56.4\%$$169_{\pm.017}.$$405_{\pm.079}.$$802_{\pm.055}.$ $45.9\%$
+SMOTE .$331_{\pm.017}.$$413_{\pm.025}.$$777_{\pm.028}.$ $50.7\%$$276_{\pm.045}.$$359_{\pm.026}.$$826_{\pm.041}.$ $48.7\%$
+STaSy .$304_{\pm.072}.$$532_{\pm.060}.$$780_{\pm.072}.$ $53.9\%$$240_{\pm.106}.$$542_{\pm.065}.$$739_{\pm.073}.$ $50.7\%$
+TabDDPM .$281_{\pm.004}.$$341_{\pm.017}.$$750_{\pm.031}.$ $45.7\%$$191_{\pm.010}.$$339_{\pm.010}.$$816_{\pm.040}.$ $44.9\%$
+TabSyn .$286_{\pm.005}.$$319_{\pm.018}.$$815_{\pm.016}.$ $47.3\%$$182_{\pm.011}.$$285_{\pm.022}.$$847_{\pm.016}.$ $43.8\%$
+FairCB .$305_{\pm.005}.$$623_{\pm.011}.$$762_{\pm.023}.$ $56.3\%$$227_{\pm.006}.$$680_{\pm.021}.$$824_{\pm.023}.$ $57.7\%$
+FairTGAN .$481_{\pm.021}.$$812_{\pm.115}.$$692_{\pm.026}.$ $66.2\%$$544_{\pm.036}.$$567_{\pm.169}.$$732_{\pm.024}.$ $61.4\%$
+Ours .$503_{\pm.017}.$$690_{\pm.085}.$$781_{\pm.006}.$ $65.8\%$$580_{\pm.027}.$$631_{\pm.087}.$$781_{\pm.015}.$ $66.4\%$